### v1

# Provable FDR Control for Deep Feature Selection: Arbitrarily Deep MLPs and Beyond

Abstract.
We develop a flexible feature selection framework based on deep neural networks that approximately controls the false discovery rate (FDR), a measure of Type-I error. The method applies to architectures whose first layer is fully connected. From the second layer onward, it accommodates multilayer perceptrons of arbitrary width and depth, convolutional and recurrent networks, attention mechanisms, residual connections, and dropout. The procedure also accommodates stochastic gradient descent with data-independent initializations and learning rates. To the best of our knowledge, this is the first work to provide a theoretical guarantee of FDR control for feature selection within such a general deep learning setting.
Our analysis is built upon a multi-index data-generating model and an asymptotic regime in which the feature dimension $n$ diverges faster than the latent dimension $q^{*}$, while the sample size, the number of training iterations, the network depth, and hidden layer widths are left unrestricted. Under this setting, we show that each coordinate of the gradient-based feature-importance vector admits a marginal normal approximation, thereby supporting the validity of asymptotic FDR control. As a theoretical limitation, we assume $\bm{B}$-right orthogonal invariance of the design matrix, and we discuss broader generalizations. We also present numerical experiments that underscore the theoretical findings.
Contact: sawaya@g.ecc.u-tokyo.ac.jp
KS is supported by JST ACT-X (JPMJAX24CC) and Grant-in-Aid for JSPS Fellows (24KJ0841).
1. Introduction
Feature selection is the task of identifying features that are truly relevant to the response $\bm{y}$. It plays a dual role in modern machine learning: it underpins scientific discovery, such as identifying genes associated with Alzheimer’s disease, and it enhances the interpretability of predictive models. Two largely separate research threads have pursued this goal, namely high-dimensional statistics and explainable AI (XAI).
On the statistics side, a central objective has been to provide theoretical guarantees on false discovery rate (FDR) control [abramovich2006adapting, meinshausen2006high, barber2015controlling, candes2018panning, xing2023controlling, dai2023false, dai2023scale, du2023false], often by debiasing sparse estimators like the LASSO [tibshirani1996regression]. These guarantees, albeit rigorous, typically come at the expense of strong modeling assumptions (e.g., the true model follows a generalized linear model, or the feature distribution is known), which can diverge from complex real‑world phenomena and thereby cast doubt on the reliability of the selected features.
On the XAI side, a wealth of attribution methods, e.g., LIME [ribeiro2016should], SHAP [lundberg2017unified], random forest feature importance [breiman2001random], and saliency maps [simonyan2013deep], quantify the contribution of individual features in black‑box models. Thresholding such scores yields a practical selection heuristic, yet without guarantees on Type‑I error; thus, error control has remained elusive.
This paper bridges these threads. We propose a feature‑selection procedure for deep neural networks that approximately controls the FDR while retaining the modeling flexibility of modern architectures. Our approach is built upon input sensitivity $\bm{\xi}^{(t)}\in\mathbb{R}^{n}$ defined by the gradient of the trained neural network’s output with respect to input, and a simple data‑splitting aggregation scheme. The analysis sheds light on when input sensitivity admits a normal approximation and how this leads to valid error control, irrespective of architecture details. Our main contributions are:
  • Flexible scope across architectures. We develop a feature-selection method applicable to multilayer perceptrons (MLPs) with arbitrary width and depth, as well as convolutional and recurrent networks, attention mechanisms, residual connections, and dropout.
  • Agnostic to the training protocol. Our guarantees accommodate stochastic gradient descent with arbitrary, data-independent initialization schemes and learning rates.
  • Normal approximation of input sensitivity. We show that input sensitivity $\bm{\xi}^{(t)}$ is asymptotically normal when the feature dimension $n$ is sufficiently larger than the latent dimension $q^{*}$. This holds regardless of the sample size $m$ and all the network parameters (e.g., width and depth). Also, this result holds at each training iteration $t$, enabling early stopping.
  • Asymptotic FDR control via sample splitting. By aggregating input sensitivity across splits, our procedure achieves asymptotic FDR control with a simple and implementable pipeline.
  • General data‑generating process. The theory is established under a multi‑index model as the data-generating process with unknown nonlinearity. This is a flexible framework that captures rich latent structures beyond generalized linear models.
  • New proof technique. At the crux of the analysis is a technique relying on the recursive inheritance of orthogonal invariance of the input sensitivity, which may be of independent interest.
  • Empirical support. Numerical experiments underscore that the normal approximation and FDR control hold under the stated conditions, aligning with the theory.
On the other hand, a major limitation of our framework lies in the assumption that the design matrix $\bm{X}\in\mathbb{R}^{m\times n}$ is $\bm{B}$-right orthogonally invariant (See Assumption 1 (ii) for the definition). While this assumption accommodates time dependence, heavy-tailed distributions, and low-rank structures, it does not allow us to account for specific forms of feature correlations (See Appendix B for details). To address this limitation, we discuss potential extensions to the general correlation structures in Appendix C.
It should be noted that FDR control represents only the minimal requirement of avoiding excessive inclusion of irrelevant variables in feature selection. With respect to the complementary criterion of Type-II error (the ability to correctly identify truly relevant features), our analysis, like much of the existing literature, provides no theoretical guarantees, although numerical evaluations are reported. For instance, a procedure that selects nothing trivially attains an FDR of zero, yet suffers a Type-II error of one. A comprehensive assessment of feature selection methods therefore requires attention to both criteria.
1.1. Related works
1.1.1. FDR control in feature selection.
There are two main research streams of FDR control in feature selection: estimator-based FDR control and knockoff filters. In the former, e.g., for linear regression, whether coefficients are zero or nonzero determines which features should be selected, and FDR control can be achieved by invoking asymptotic normality of the estimators. Representative examples include the Gaussian mirror [xing2023controlling] and data splitting [dai2023false]. A key advantage of these approaches is scale-free property, since they avoid estimating the asymptotic variance, and they have been extended to generalized linear models [dai2023scale]. Similar techniques have also been applied to sliced inverse regression under multi-index models [zhao2022testing], although the theoretical assumptions on the feature dimension and the sparsity level are rather stringent.
As the second line of research, the model-X knockoff [candes2018panning] is groundbreaking in that it can control the FDR without assuming the structure of $\bm{y}\mid\bm{X}$ where $\bm{y}\in\mathbb{R}^{m}$ is a response vector. However, it requires knowledge of the joint distribution of $\bm{X}$, which is restrictive. To address this limitation, methods that estimate the distribution of $\bm{X}$ using generative networks have been proposed [jordon2019knockoffgan, romano2020deep]. In addition, a sequential knockoff sampler for a given feature distribution has been proposed [bates2021metropolized].
1.1.2. FDR control via neural networks.
Several studies have proposed FDR control methods that employ neural networks, although without theoretical guarantees and typically within restricted classes of architectures. The Neural Gaussian Mirror [xin2020ngm] defines a kernel-based conditional dependence measure and performs feature selection with MLPs. DeepPINK [lu2018deeppink] is a knockoff framework with a specially designed network architecture in which the features are assumed to be jointly Gaussian, and DeepLINK [zhu2021deeplink] relaxes this distributional restriction. These approaches have been shown empirically to achieve FDR control. Nevertheless, because theoretical guarantees are not provided, it remains unclear under what conditions FDR control is achievable, and the range of supported network architectures is limited.
1.2. Notations
Vectors and matrices are typeset in boldface (e.g., $\bm{x},\bm{B}$). For $n\in\mathbb{N}$, $n]=\{1,\ldots,n\}$. For $S\subset[n]$, $S^{\mathsf{c}}=[n]\setminus S$. For $\bm{a}\in\mathbb{R}^{n}$ and $S\subset[n]$, we denote by $\bm{a}_{S}$ the subvector of $\bm{a}$ consisting of the entries indexed by $S$. For a matrix $\bm{A}$, let $\bm{A}^{+}$ be the Moore-Penrose pseudo-inverse of $\bm{A}$. $\Phi:\mathbb{R}\to\mathbb{R}$ and $\phi:\mathbb{R}\to\mathbb{R}$ are the cumulative distribution function and the density function of the standard Gaussian distribution, respectively.
2. Problem formulation
Suppose that we observe a response vector $\bm{y}=(y_{1},\ldots,y_{m})^{\top}\in\mathbb{R}^{m}$ together with a design matrix $\bm{X}=(\bm{x}_{1},\ldots,\bm{x}_{m})^{\top}\in\mathbb{R}^{m\times n}$, where $m$ is the sample size and $n$ is the number of features. Our goal is to select a relevant subset of feature indices from $[n]$ that are associated with $\bm{y}$. A desirable feature selection procedure controls the false discovery rate (FDR) [benjamini1995controlling], defined by
$\textsf{FDR}=\mathbb{E}[\textsf{FDP}],\quad\textsf{FDP}=\frac{\#\{j\notin S:j\in\hat{S}\}}{\#\{j:j\in\hat{S}\}\vee 1},$ (1)
where $S$ denotes the set of indices corresponding to the relevant features, and $\hat{S}$ is the set of selected indices.
We consider the situation where $(\bm{y},\bm{X})$ follows the multi-index model defined below.
Definition 1 (Multi-index model).
We say that a pair of the response vector $\bm{y}\in\mathbb{R}^{m}$ and the design matrix $\bm{X}\in\mathbb{R}^{m\times n}$ follows a multi-index model if there exists a weight matrix $\bm{B}\in\mathbb{R}^{n\times q^{*}}$, a deterministic function $g:\mathbb{R}^{q^{*}}\times\mathbb{R}\to\mathbb{R}$, and noise variables $\bm{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{m})^{\top}$ independent of $\bm{X}$ such that, for each $i\in[m]$,
$y_{i}=g(\bm{B}^{\top}\bm{x}_{i},\varepsilon_{i}).$ (2)
Let $\bm{B}=(\bm{b}_{1},\ldots,\bm{b}_{n})^{\top}$. This formulation (2) encompasses linear regression, logistic regression, and certain neural network models with $q^{*}$ hidden units in the first layer. The column space of $\bm{B}$ is often referred to as the central subspace of $\bm{X}$.
We now consider fitting a neural network to the observations. Let $f_{\mathcal{W}}:\mathbb{R}^{n}\to\mathbb{R}$ denote a neural network parameterized by the set $\mathcal{W}$, e.g., including weight matrices $(\bm{W}_{1},\ldots,\bm{W}_{L})$, where $L$ is the number of layers. Given a loss function $\mathcal{L}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, the empirical risk minimization problem is
$\min_{\mathcal{W}}\sum_{i=1}^{m}\mathcal{L}(y_{i},f_{\mathcal{W}}(\bm{x}_{i})).$ (3)
For example, we may use the quadratic loss $\mathcal{L}(u,v)=(u-v)^{2}$ for regression and the cross-entropy loss $\mathcal{L}(u,v)=\log(1+\exp(v))-uv$ for binary classification. The optimization is performed by (stochastic) gradient descent, starting from initial parameters $\mathcal{W}^{(0)}$ and yielding updated parameters $\mathcal{W}^{(t)}$ after $t$ iterations.
After training, we evaluate feature importance by the partial derivative of the fitted network with respect to each input feature. Specifically, for $t\in\mathbb{N}$ and $j\in[n]$, define
$\xi_{j}^{(t)}\equiv\sum_{i=1}^{m}\frac{\partial}{\partial x_{ij}}f_{\mathcal{W}^{(t)}}(\bm{x}_{i}).$ (4)
If the fitted network is differentiable almost everywhere, the input sensitivity $\xi_{j}^{(t)}$ can serve as a measure of the contribution of the $j$-th feature to the response $\bm{y}$. After computing $\bm{\xi}^{(t)}=(\xi_{1}^{(t)},\ldots,\xi_{n}^{(t)})^{\top}$, we then determine an appropriate cutoff to control the FDR.
3. Theoretical background
To control the FDR, we need marginal distributional characterizations of the input sensitivity $\bm{\xi}^{(t)}$ under the null. In this section, we characterize the distribution of suitably transformed $\bm{\xi}^{(t)}$ for each $n,m,q^{*},t\in\mathbb{N}$. Based on this, we establish the marginal asymptotic normality of the feature importance uniformly under the null as $n\to\infty$ with $q^{*}=o(n)$, for arbitrary $m$ and $t$.
In what follows, we formally define the index set $S^{\mathsf{c}}$ of null features.
Definition 2.
We say that $x_{j}$ for $j\in[n]$ is a null feature if,
$y\perp\!\!\!\perp x_{j}\mid\bm{x}_{-j}.$ (5)
We then define $S^{\mathsf{c}}$ as the index set of all null features.
Assumption 1 (Multi-index model and $\bm{B}$-ROI design).
(i) The observation $(\bm{y},\bm{X})\in\mathbb{R}^{m}\times\mathbb{R}^{m\times n}$ follows the multi-index model in Definition 1 with a full rank $\bm{B}\in\mathbb{R}^{n\times q^{*}}$.
(ii) The design matrix $\bm{X}\in\mathbb{R}^{m\times n}$ satisfies $\bm{X}\overset{\rm d}{=}\bm{X}\bm{U}$ for any orthogonal matrix $\bm{U}\in\mathbb{R}^{n\times n}$ such that $\bm{U}\bm{B}=\bm{B}$. We shall call this property $\bm{B}$-ROI in the sequel.
The design assumption (ii) permits row-wise dependence, heavy-tailed marginals, and low-rank structures. Still, it rules out certain forms of column dependence, discrete-valued entries, and multi-modal distributions. See Appendix B for further discussion. We also discuss the robustness to elliptical designs in Section C. Additional mild regularity conditions under which (5) holds if and only if $\bm{b}_{j}=\bm{0}_{q^{*}}$ are provided in Appendix A.4.
Assumption 2 (Loss function).
Suppose that the loss function $\mathcal{L}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ has a finite partial derivative with respect to the second argument ${\partial_{2}}\mathcal{L}(u,v)$ for almost every $v\in\mathbb{R}$.
This covers most losses encountered in practice.
Assumption 3 (Architecture of the neural network).
The first layer of the network is taken to be dense and fully connected, and we denote its weight matrix by $\bm{W}_{1}\in\mathbb{R}^{n\times q}$. The dependence of the entire network $f_{\mathcal{W}}(\bm{x})$ on any input $\bm{x}\in\mathbb{R}^{n}$ arises solely through the transformed representation $\bm{W}_{1}^{\top}\bm{x}$.
This assumption still encompasses multilayer perceptrons with arbitrary width, depth, and activation functions. From the second layer onward, we allow any structures, including residual connections and dropout, which is a pragmatic modeling choice.
Intuitively, the linear representation $\bm{W}_{1}^{\top}\bm{x}$ serves as a surrogate for $\bm{B}^{\top}\bm{x}$ in the multi-index model, while the network’s subsequent nonlinearity approximates $g(\cdot)$. The matrix sizes of $\bm{W}_{1}$ and $\bm{B}$ need not match.
Assumption 4 (SGD options).
(i) Every element of initial parameters $\mathcal{W}^{(0)}$ is independent of $(\bm{y},\bm{X})$, $\bm{W}_{1}^{(0)}$ and $\mathcal{W}_{\setminus 1}^{(0)}$ are independent, and $\bm{W}_{1}^{(0)}$ satisfies $\tilde{\bm{U}}\bm{W}_{1}^{(0)}\overset{\rm d}{=}\bm{W}_{1}^{(0)}$ for any orthogonal matrix $\tilde{\bm{U}}\in\mathbb{R}^{n\times n}$.
(ii) Let the mini-batch indices $I_{t}\subseteq[m]$ and the learning rate $\eta_{t}>0$ be independent of $(\bm{y},\bm{X},\mathcal{W}^{(0)})$ for all $t\in\mathbb{N}$.
For instance, the entrywise i.i.d. Gaussian $\bm{W}_{1}^{(0)}$ with any common variance, including He- and Xavier-initializations, satisfies the assumption.
Under these assumptions, we obtain the following.
Proposition 1.
Under Assumptions 14, for each $m,n,q^{*}\in\mathbb{N}$ and iteration $t\in\mathbb{N}$ of the SGD with/without replacement, conditioning on the learning-rate and mini-batch schedule,
$\frac{\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}$ (6)
is uniformly distributed on the unit sphere lying in $\mathrm{Col}(\bm{B})^{\perp}$. Here, $\mathrm{Col}(\bm{B})^{\perp}$ is the orthogonal complement of the column space of $\bm{B}$, and $\bm{P}_{\bm{B}}^{\perp}=\bm{I}_{n}-\bm{B}(\bm{B}^{\top}\bm{B})^{+}\bm{B}^{\top}$ is the orthogonal projection matrix onto $\mathrm{Col}(\bm{B})^{\perp}$.
Figure 1 provides an illustration of Proposition 1. It shows that ${\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}$ is uniformly distributed around $\bm{B}$ while maintaining a constant angle. Since $\bm{B}$ itself reflects the intrinsic importance of each feature, this observation supports the consistency of interpreting $\bm{\xi}^{(t)}$ as feature importance.
Figure 1. A schematic illustration of $\bm{\xi}^{(t)}$ for $q^{*}=1$ and $n=3$. $\bm{U}\in\mathbb{R}^{n\times n}$ is any orthogonal matrix such that $\bm{U}\bm{B}=\bm{B}$ (i.e., rotation around $\bm{B}$).
The proof is completed by replacing the orthogonal invariance to be established for $\bm{\xi}^{(t)}$ with respect to $\bm{B}$ by an equivalent invariance of the first-layer weights $\bm{W}_{1}^{(t)}$ via the chain rule, and then showing recursively in $t$ that this invariance is preserved by the update.
From spherical uniformity it follows that, letting $\bm{Z}\sim\mathcal{N}(\mathbf{0},\bm{I}_{n})$,
$\frac{\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}\ \overset{\mathrm{d}}{=}\ \frac{\bm{P}_{\bm{B}}^{\perp}\bm{Z}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{Z}\|}.$
Together with the fact that for each null feature index $j\in S^{\mathrm{c}}$ we have $\bm{b}_{j}=\mathbf{0}_{q^{\star}}$, this yields the following asymptotic normality:
Theorem 1.
Under Assumptions 14, for any null feature index $j\in S^{\mathsf{c}}$, we have
$\frac{\sqrt{n}\xi_{j}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}\overset{\mathrm{d}}{\to}\mathcal{N}(0,1),$ (7)
as $n\to\infty$ while $q^{*}=o(n)$. Furthermore, the convergence holds uniformly in $j\in S^{\mathsf{c}}$ in the sense that
$\sup_{j\in S^{\mathsf{c}}}\ \sup_{u\in\mathbb{R}}\left\lvert\,\mathbb{P}\left(\frac{\sqrt{n}\,\xi_{j}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}\leq u\right)-\Phi(u)\,\right\rvert\ \to\ 0.$ (8)
From this theorem, the asymptotic null distribution of ${\xi_{j}^{(t)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}$ is identified. However, estimation of $\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|$ is challenging since it depends on the unknown structure $\bm{B}$. In the next section, we demonstrate that by employing the data-splitting technique, the multiplicative factor that appears uniformly across all $j\in[n]$ can be ignored (the scale-free property), which enables valid FDR control.
Proofs of the assertions in this section are deferred to Appendix A. Figure 2 exhibits that numerical results confirm the asymptotic normality of Theorem 1.
The technical contribution underlying Proposition 1 and Theorem 1 is to extend the orthogonal-invariance-based theory of marginal asymptotic normality—originally developed by zhao2022asymptotic and later shown to be broadly applicable by sawaya2024high—from loss minimizers (i.e., M-estimators) to the individual iterates of loss-minimization algorithms. As a consequence, our asymptotic normality results do not rely on
  1. (i)
    the existence or uniqueness of a loss minimizer,
  2. (ii)
    convexity of the loss function,
  3. (iii)
    or convergence of the optimization path.
The availability of early stopping is also appealing, as it may help reduce type-II error. Points (ii) and (iii) are particularly important in the training of deep neural networks, where convergence to a global minimum is rarely guaranteed. Furthermore, our proofs extend naturally to gradient-based procedures for fitting conventional M-estimators without neural networks. This implies that valid inference remains possible even in settings where a maximum likelihood estimator may fail to exist—as in logistic regression [candes2020phase]—which further underscores the practical value of our approach.
4. Methodology
In this section, we construct the actual feature selection procedure. It consists of two stages: (I) computing an importance statistic $M_{j}$ for each feature that possesses desirable distributional properties, and (II) selecting features by applying an appropriate thresholding rule that controls the FDR. The properties required for the statistic to be scale-free in (I) are as follows:
  • (a)
    If $j\in S^{\mathsf{c}}$, then $M_{j}$ follows (asymptotically) a distribution symmetric around zero.
  • (b)
    If $j\in S$, then $M_{j}$ takes large positive values.
According to Theorem 1, $\xi_{j}^{(t)}$ satisfies (a) but not (b), whereas $|\xi_{j}^{(t)}|$ satisfies (b) but not (a). Therefore, we adopt a data-splitting approach [dai2023false, dai2023scale]. Specifically, we randomly divide the data into two equal parts and, using the quantities $\xi_{j1}^{(t)}$ and $\xi_{j2}^{(t)}$ computed from each split, construct the importance statistics, for $j\in[n]$,
$M_{j}=\mathsf{sign}\left(\xi_{j1}^{(t)}\xi_{j2}^{(t)}\right)\psi\left(|\xi_{j1}^{(t)}|,\,|\xi_{j2}^{(t)}|\right),$
where $\psi:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is a user-specified function assumed to be non-negative, symmetric, positive homogeneous, and monotone in each input, e.g., $\psi(u,v)=uv$, $\min(u,v)$, and $u+v$.
The crucial point here is that, unlike p-value based methods for FDR control, which require knowledge of the entire null distribution, this approach only relies on the symmetry under the null. As a result, no information about the asymptotic variance or convergence rate of the limiting null distribution of $M_{j}$ is required.
The intuition behind why symmetry alone suffices is as follows. We can determine the cutoff for the nominal level $\alpha\in(0,1)$ as
$\tau_{\alpha}=\min\left\{u>0:\widehat{\mathsf{FDP}}(u)\equiv\frac{\#\{j:M_{j}<-u\}}{\#\{j:M_{j}>u\}\vee 1}\leq\alpha\right\}.$ (9)
This is expected to control the FDR because, if $M_{j}$ is symmetric around zero under the null, we have
$\mathsf{FDP}(u)$ $=\frac{\#\{j\in S^{\mathsf{c}}:M_{j}>u\}}{\#\{j:M_{j}>u\}\vee 1}\overset{\rm d}{=}\frac{\#\{j\in S^{\mathsf{c}}:M_{j}<-u\}}{\#\{j:M_{j}>u\}\vee 1}$ (10)
$\leq\frac{\#\{j:M_{j}<-u\}}{\#\{j:M_{j}>u\}\vee 1}=\widehat{\mathsf{FDP}}(u).$ (11)
Overall procedure is summarized in Algorithm 1.
0: Nominal level $\alpha\in(0,1)$, the observation $(\bm{y},\bm{X})\in\mathbb{R}^{m}\times\mathbb{R}^{m\times n}$, the stopping time $T\in\mathbb{N}$, and $\psi:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$.
1: Split the data into two equal-sized halves $(\bm{X}^{(1)},\bm{y}^{(1)})$ and $(\bm{X}^{(2)},\bm{y}^{(2)})$.
2: For each part of the data, calculate $\bm{\xi}_{\cdot 1}^{(T)}$ and $\bm{\xi}_{\cdot 2}^{(T)}$ as in (4) after $T$ updates of the SGD.
3: Obtain the importance statistics $M_{j}=\mathsf{sign}\left(\xi_{j1}^{(T)}\xi_{j2}^{(T)}\right)\,\psi\left(|\xi_{j1}^{(T)}|,\,|\xi_{j2}^{(T)}|\right)$ for each $j\in[n]$.
4: Select features above the cutoff $\tau_{\alpha}=\min\{u>0:\widehat{\mathsf{FDP}}(u)\leq\alpha\}$.
Algorithm 1 pseudocode for the selection procedure
This selection procedure asymptotically controls the FDR at a predetermined level under additional assumptions.
Assumption 5.
$\psi(\cdot,\cdot)$ is non-negative, symmetric about two inputs, and monotone in each input. Additionally, there exists $r>0$ such that for all $a\geq 0$ and $(s,t)\in[0,\infty)^{2}$,
$\psi(as,at)\;=\;a^{r}\,\psi(s,t).$
This is the formal requirement imposed on the user-specified function $\psi$.
Assumption 6.
In Algorithm 1, suppose that $(\bm{X}^{(1)},\bm{y}^{(1)})\overset{\rm d}{=}(\bm{X}^{(2)},\bm{y}^{(2)})$. Additionally, assume that the construction of $\bm{\xi}_{\cdot 1}^{(T)}$ and $\bm{\xi}_{\cdot 2}^{(T)}$ is the same; for example, the randomness of initializations, learning rate, and loss function are common.
This assumption is necessary for the validity of data splitting.
Assumption 7.
Let $S^{+}(u)=\#\{j\in S:M_{j}>u\}$, $S^{-}(u)=\#\{j\in S:M_{j}<-u\}$, and $S^{\pm}(u)=S^{+}(u)+S^{-}(u)$. There exist $c,\theta,\rho\in(0,1)$ such that, for $K_{n}=\lfloor cn\rfloor$, as $n\to\infty$,
$\mathbb{P}\left(S^{\pm}(u_{K_{n}})\geq\theta K_{n},~~\inf_{0\leq u\leq u_{K_{n}}}\frac{S^{+}(u)}{S^{\pm}(u)}\geq\rho\right)\to 1,$ (12)
where $u_{K_{n}}$ is the $K_{n}$-th largest magnitude among $M_{j}$’s. Moreover, the following holds with the given nominal level $\alpha$:
$(\alpha\rho-(1-\rho))\,\theta\;>\;\frac{1-\alpha}{2}\,(1-\theta).$ (13)
Such assumptions frequently appear in the related literature. Compared with Assumption 3.2 in dai2023scale, which requires that a fixed proportion of the true signals diverge, our assumption can be regarded as considerably weaker.
Assumption 7 requires that, within the top $K_{n}$ statistics $M_{j}$’s ranked by magnitude, at least a fixed fraction corresponds to non-null and, moreover, lies on the positive side. That is, among the non-null variables, at least a certain fraction is required to possess genuinely positive importance scores, and this requirement is expected to hold increasingly as the iteration $t$ advances. Denote $n_{0}=|S^{\mathsf{c}}|$.
Theorem 2.
Suppose Assumptions 17 hold and $n_{0}/n\to\pi_{0}\in(0,1]$. Then, Algorithm 1 satisfies, for any nominal level $\alpha\in(0,1)$,
$\mathsf{FDP}\leq\alpha+o_{\mathsf{p}}(1)\quad\mathrm{and}\quad\limsup_{n\to\infty}\,\mathsf{FDR}\leq\alpha.$ (14)
A drawback of the data-splitting method is that it effectively halves the sample size, which may reduce power. One remedy, known in the literature and also applicable here, is to aggregate the selection results obtained from multiple random splits of the data [dai2023false]. Related stabilization techniques include du2023false and ren2024derandomised
5. Numerical experiments
In this section, we empirically validate the theoretical guarantees developed in the preceding sections. We then compare the proposed method against relevant baselines. All code and scripts for reproducing our results is available at https://github.com/sawaya-ka/deep-feature-selection. Further experiments are provided in Appendix D.
5.1. Marginal asymptotic normality
We numerically verify Theorem 1. The data are generated according to
$y=g(\bm{b}_{1}^{\top}\bm{x})+\sum_{k=2}^{q^{*}}\left\{h(\bm{b}_{k}^{\top}\bm{x})\cdot(\bm{b}_{k-1}^{\top}\bm{x})\right\}+\varepsilon,$ (15)
with $\varepsilon\sim\mathcal{N}(0,1)$, $\bm{x}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$, $q^{*}=8$, $g(x)=(x-2)^{2}$ and $h(x)=\max(x,0)$. The vector $\bm{b}_{1}$ has its first half of entries equal to $2/\sqrt{n}$ and the remaining entries are zero, and $\bm{b}_{k}=\bm{e}_{k}\in\mathbb{R}^{n}$ for $k=2,\ldots,q^{*}$.
We fixed the batch size of SGD to 128 and the learning rate to $3\times 10^{-3}$ except for Transformer. After ten update steps, we constructed a histogram of the latter half of the components of ${\sqrt{n}\bm{\xi}^{(10)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(10)}\|}$ and compared it with the density of a normal distribution whose mean and variance match the sample mean and sample variance of these components, as well as with the standard normal density $\mathcal{N}(0,1)$. The resulting plots are presented in Figure 2.
Figure 2. Histograms of the empirical distribution of ${\sqrt{n}\xi_{j}^{(10)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(10)}\|}$ for $j\in S^{\mathsf{c}}$. The solid black curve shows the $\mathcal{N}(0,1)$ density. The solid red curve represents a normal density fitted to the histograms, and the dotted blue line indicates the empirical mean.
Also, the corresponding $Q$-plots are provided in Figure 6 of Appendix D.
The settings labeled $n\ll m$, $n\asymp m$, and $n\gg m$ correspond respectively to $(m,n)=(100000,1000)$, $(2000,1000)$, and $(10,1000)$. The network architectures used for training in these experiments are described below. We employ dropout with a rate $0.1$.
5-layer MLP. We use a 5-layer MLP consisting of four hidden layers of widths $(1024,1024,512,256)$ with ReLU activations and a final linear output. All weights are initialized with He-normal initialization.
1D-CNN. The 1D-CNN baseline first applies the lifting layer $\bm{W}_{1}^{\top}\bm{x}$ with $q=n$ followed by a stack of three convolutional layers with output channels $(64,128,128)$, kernel sizes $(11,9,7)$, and stride $1$. Dilated convolutions with dilation rates $(1,2,4)$ are used to enlarge the receptive field. ReLU activation is applied after each convolution. The convolutional output is flattened and passed through a fully connected head with hidden layers $(128,64)$ and a final linear output.
LSTM. For sequential modeling we adopt a two-layer bidirectional LSTM with hidden size $128$. Each time step input is obtained by projecting the lifted representation $\bm{W}_{1}^{\top}\bm{x}$ into dimension $4$. The last hidden states of the forward and backward directions are concatenated and fed into a fully connected head with hidden layer $(64)$ and linear output. We use Xavier-normal initialization for the recurrent weights.
Transformer. The input is first lifted by a trainable dense map $\bm{W}_{1}^{\top}\in\mathbb{R}^{q\times n}$ (Xavier initialization), producing $q$ tokens. Each token is embedded by a linear ID-specific map and normalized (LayerNorm), without positional encoding. We employ a Transformer encoder with two layers, model dimension $256$, four heads, feed-forward dimension $256$, and GELU activation. For sequence aggregation we use gated pooling with hidden dimension $32$ and temperature $\tau=1.0$, followed by a fully connected head with hidden size $32$ and linear output. Training uses AdamW with $(\beta_{1},\beta_{2},\varepsilon)=(0.9,0.95,10^{-8})$, weight decay $0.01$, batch size $128$, base learning rate $3\!\times\!10^{-4}$ with $10\%$ warm-up, and gradient clipping at $1.0$.
These experiments were conducted on a Google Cloud Platform VM equipped with a single NVIDIA A100 (40GB) GPU using PyTorch 2.6.0 with CUDA 12.4.
5.2. FDR control
We next demonstrate that Algorithm 1 is able to approximately control the false discovery rate (FDR) at or below the nominal level $\alpha=0.1$ under the stated assumptions. We consider the setting $(m,n)=(1600,400)$; the data-generating process and learning architectures are otherwise the same as in the previous section. We use $\psi(u,v)=\min(u,v)$.
We define the power in the feature selection problem:
$\mathsf{Power}=\mathbb{E}\left[\frac{\#\{j\in S:j\in\hat{S}\}}{\#\{j:j\in S\}}\right].$ (16)
We can see that the power is one minus the Type-II error.
Figure 3 reports the results. As anticipated in Assumption 7, once training progresses and the power reaches a reasonable level, the FDR is also brought under control. The trajectory of the training loss corresponding to these experiments is shown in Figure 7 of Appendix D.
Experiments were conducted on a Google Cloud Platform VM equipped with four NVIDIA Tesla T4 (16 GB each) GPUs, using PyTorch 2.8.0 (built with CUDA 12.8) and CUDA runtime 12.4.
Figure 3. Results for the false discovery rate (left) and power (right) when performing feature selection at each iteration using the artificial data and models defined in Section 5.1. The solid curves represent averages over 20 independent runs, and the shaded areas indicate one standard deviation around the mean.
Remark 5.1.
In a single-index model $y=g(\langle\bm{w}_{\star},\bm{x}\rangle)+\varepsilon$, if the trained network well approximates the regression function then $\nabla_{\bm{x}}f_{\mathcal{W}^{(t)}}(\bm{x})\approx g^{\prime}(\langle\bm{w}_{\star},\bm{x}\rangle)\,\bm{w}_{\star}$, so the averaged gradient $\bm{\xi}^{(t)}$ converges to $\mathbb{E}[g^{\prime}(Z)]\,\bm{w}_{\star}$ with $Z\sim\mathcal{N}(0,\|\bm{w}_{\star}\|^{2})$. While the factor $\mathbb{E}[g^{\prime}(Z)]$ being close to zero does not affect the validity of FDR control, it cancels out the signal direction and thereby reduces the power of Algorithm 1.
5.3. Comparison with other methods
5.3.1. Competing methods.
We benchmark our approach against flexible variable-selection baselines that can (or aim to) control FDR. Neural baselines advertised as enabling FDR control via deep representations include the Neural Gaussian Mirror (NGM) [xin2020ngm] and DeepLINK [zhu2021deeplink]. While empirically competitive, these methods do not offer formal guarantees. All methods are run at the same nominal FDR level $\alpha=0.1$.
We ran DeepLINK [zhu2021deeplink] with authors’ official code [^1] https://github.com/zifanzhu/DeepLINK with hyperparameters as follows: L1 penalty $10^{-3}$, learning rate $10^{-3}$, ELU activation, and mean squared error loss. Column-wise centering and scaling were applied by default.
Among the methods compared, NGM was thus far the most computationally demanding. To make large-scale experiments feasible, we employed several approximations: a preliminary screening step retaining at most $n/2$ variables, a subsample-based approximation of the kernel matrices (subsample size $800$), and a coarse grid of 6 candidate values for the scale parameter $c_{j}$. The neural network used in NGM was a two-hidden-layer MLP with hidden widths proportional to $\log n$, trained for 60 epochs with batch size 256 and learning rate $10^{-3}$.
5.3.2. Data generating process.
We fix the ambient dimension at $n=500$ and vary the sample size $m\in\{2000,\,1000,\,500\}$ to probe the effect of sample scarcity. Unless otherwise noted, the data-generating process follows Section 5.1 exactly, except that (because the design $\bm{X}$ is scaled by $1/\sqrt{n}$) we set the nonzero entries of the signal matrix $\bm{B}$ to $2$ (rather than $2/\sqrt{n}$).
We consider four scenarios for the design matrix $\bm{X}\in\mathbb{R}^{m\times n}$ satisfying Assumption 1 (ii).
N(0,1). Entries are i.i.d. Gaussian: $X_{ij}\overset{\rm iid}{\sim}\mathcal{N}(0,1/n),\,(i,j)\in[m]\times[n]$.
t(3). Entries are i.i.d. standardized $t$: $X_{ij}\overset{\rm iid}{\sim}\tfrac{1}{\sqrt{n}}t(3),\,(i,j)\in[m]\times[n],$ where $t(3)$ denotes the $t$-distribution with $3$ degrees of freedom.
Spiked. Let $r=2$ be the spike rank. We write $\bm{X}=\bm{V}_{1}\bm{V}_{2}^{\top}+\bm{E},$ where $\bm{V}_{1}\in\mathbb{R}^{m\times r}$ and $\bm{V}_{2}\in\mathbb{R}^{n\times r}$ have orthonormal columns (each drawn as $r$-columns from independent Haar orthogonal matrices of sizes $m$ and $n$, respectively), and $\bm{E}$ has i.i.d. entries $E_{ij}\sim\mathcal{N}(0,1/n)$.
Our inferential procedures employ sample splitting: we partition the $m$ rows into two equal halves, $\bm{X}^{(A)}\in\mathbb{R}^{(m/2)\times n}$ and $\bm{X}^{(B)}\in\mathbb{R}^{(m/2)\times n}$. For N(0,1) and t(3), closure under row-subsampling is immediate. For Spiked, we assume a modeling convention such that, when the $m$ rows are partitioned into two equal halves, each half independently follows the same distributional form as the original model (with common latent parameters), so that both $\bm{X}^{(A)}$ and $\bm{X}^{(B)}$ are valid Spiked samples.
Table 1. Performance comparison for $m=2000$.
MLP 1D-CNN LSTM DeepLINK NGM
Design FDR Power FDR Power FDR Power FDR Power FDR Power
N(0,1) 0.075 (0.019) 0.830 (0.049) 0.094 (0.028) 0.998 (0.002) 0.101 (0.025) 1.000 (0.000) 0.089 (0.021) 0.999 (0.000) 0.086 (0.015) 0.851 (0.013)
t(3) 0.043 (0.042) 0.363 (0.305) 0.047 (0.044) 0.512 (0.382) 0.072 (0.040) 0.843 (0.343) 0.081 (0.030) 0.870 (0.218) 0.063 (0.031) 0.599 (0.151)
Spiked 0.076 (0.034) 0.808 (0.070) 0.099 (0.028) 0.998 (0.003) 0.104 (0.026) 1.000 (0.000) 0.087 (0.021) 1.000 (0.000) 0.080 (0.016) 0.847 (0.017)
Table 2. Performance comparison for $m=1000$.
MLP 1D-CNN LSTM DeepLINK NGM
Design FDR Power FDR Power FDR Power FDR Power FDR Power
N(0,1) 0.060 (0.030) 0.186 (0.069) 0.075 (0.031) 0.546 (0.093) 0.072 (0.024) 0.637 (0.089) 0.065 (0.021) 0.467 (0.122) 0.131 (0.031) 0.658 (0.041)
t(3) 0.053 (0.057) 0.080 (0.068) 0.040 (0.039) 0.122 (0.111) 0.061 (0.043) 0.168 (0.130) 0.064 (0.076) 0.085 (0.107) 0.144 (0.612) 0.401 (0.127)
Spiked 0.047 (0.036) 0.187 (0.066) 0.070 (0.023) 0.571 (0.094) 0.072 (0.021) 0.682 (0.089) 0.063 (0.028) 0.492 (0.126) 0.151 (0.031) 0.674 (0.046)
Table 3. Performance comparison for $m=500$.
MLP 1D-CNN LSTM DeepLINK NGM
Design FDR Power FDR Power FDR Power FDR Power FDR Power
N(0,1) 0.087 (0.092) 0.031 (0.029) 0.087 (0.079) 0.057 (0.033) 0.047 (0.068) 0.040 (0.034) 0.019 (0.58) 0.003 (0.010) 0.242 (0.049) 0.380 (0.093)
t(3) 0.074 (0.133) 0.012 (0.015) 0.041 (0.098) 0.009 (0.010) 0.058 (0.098) 0.016 (0.012) 0.017 (0.061) 0.004 (0.013) 0.325 (0.04) 0.428 (0.076)
Spiked 0.116 (0.223) 0.028 (0.033) 0.138 (0.214) 0.059 (0.046) 0.029 (0.059) 0.037 (0.036) 0.026 (0.065) 0.020 (0.049) 0.264 (0.039) 0.435 (0.056)
Tables 13 report the comparison of FDR and Power with $n=500$, averaged over 20 independent runs. The entries labeled MLP, 1D-CNN, and LSTM are defined in the previous section. It can be observed that NGM fails to control the FDR except in the $m=2000$ setting, while our method achieves relatively high power under FDR control. Furthermore, the values in parentheses below each entry denote the standard deviations computed over 20 independent random seeds. These results suggest that the occasional exceedance of the nominal FDR level $\alpha=0.1$ by our proposed method is likely due to random variation arising from averaging over a relatively small number of 20 trials. In contrast, the fact that NGM consistently yields FDR values above 0.1 with small standard deviations indicates a genuine failure in FDR control.
6. Discussion
Summary of Contributions. To the best of our knowledge, this paper provides the first theoretical guarantee for false discovery rate (FDR) control in feature selection with multilayer neural networks. This property is crucial for ensuring the reproducibility of scientific discoveries, and our results demonstrate that one can achieve both model flexibility and rigorous statistical reliability. In addition, the proposed implementation via simple data splitting is straightforward and easy to apply.
Limitations and Future Directions. Our analysis assumes that the first layer of the neural network is a dense fully connected transformation. While technically convenient, this limits the ability to exploit the spatial locality of image data or the sequential and positional structures inherent in natural language. Extending the methodology to capture richer latent structures in diverse data modalities remains an important challenge.
Moreover, we have defined the input sensitivity $\bm{\xi}^{(t)}$ as an average across instances. It would be interesting to investigate how this sensitivity can be characterized at the single-instance level, especially in domains such as computer vision where the set of pixels critical for classification may vary substantially across samples.
In this work, we considered the raw input gradients of a trained neural network as a feature importance. A promising direction for future work is to extend our analysis to more sophisticated feature attribution methods, including Integrated Gradients [sundararajan2017axiomatic], DeepLIFT [shrikumar2017learning], and SmoothGrad [smilkov2017smoothgrad]. Adapting our FDR-control framework to such attribution methods would potentially yield more stable and interpretable feature selection.
From a theoretical standpoint, our framework relied on the $\bm{B}$-right-orthogonal invariance of the design matrix $\bm{X}$. Exploring the behavior of our algorithm under more severe correlation structures is an appealing direction for future work, as is extending the theory to regularized training regimes.
Finally, while our results establish FDR control, an important practical question is which network architectures—in terms of depth, width, use of attention, dropout, or residual connections—yield higher power as functions of the sample size and ambient dimension. Answering this question is a promising research direction, though it will likely require substantially more theoretical effort.
Appendix A Proofs
A.1. Proof of Proposition 1
Proof of Proposition 1.
The basic idea of the proof is inspired by the proof of Proposition 2.1 in zhao2022asymptotic. Since $\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}/\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|$ has the unit norm and lies in $\mathrm{Col}(\bm{B})^{\perp}$, it is sufficient to show that, for any orthogonal matrix $\bm{U}\in\mathbb{R}^{n\times n}$ obeying $\bm{U}\bm{B}=\bm{B}$,
$\bm{U}\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\overset{\rm d}{=}\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}.$ (17)
We proceed to show its sufficient condition $\bm{U}\bm{W}_{1}^{(t)}=\bm{W}_{1}^{(t)}$ since Assumption 3 implies that the input sensitivity is given by, with $\mathcal{W}_{\setminus 1}^{(t)}\equiv\mathcal{W}^{(t)}\setminus\bm{W}_{1}^{(t)}$,
$\bm{\xi}^{(t)}=\sum_{i=1}^{m}\bm{W}_{1}^{(t)}h\left((\bm{W}_{1}^{(t)})^{\top}\bm{x}_{i};\mathcal{W}_{\setminus 1}^{(t)}\right),$ (18)
where $h(\cdot)$ depends on $\bm{x}_{i}$ only through $(\bm{W}_{1}^{(t)})^{\top}\bm{x}_{i}$, and given by
$h\left((\bm{W}_{1}^{(t)})^{\top}\bm{x}_{i};\mathcal{W}_{\setminus 1}^{(t)}\right)=\frac{\partial f_{\mathcal{W}^{(t)}}(\bm{x}_{i})}{\partial((\bm{W}_{1}^{(t)})^{\top}\bm{x}_{i})}.$ (19)
For each $t\in\mathbb{N}$, denote $\bm{W}_{1}^{(t)}=\bm{W}_{1}^{(t)}(\bm{y},\bm{X},\bm{W}_{1}^{(t-1)},\mathcal{W}_{\setminus 1}^{(t-1)})$ to clarify the dependence of SGD iterates on the sample $(\bm{y},\bm{X})$ and the previous iterate $\mathcal{W}^{(t-1)}=(\bm{W}_{1}^{(t-1)},\mathcal{W}_{\setminus 1}^{(t-1)})$. Then, Assumptions 2 and 3 yield, for each $t\in\mathbb{N}$,
$\bm{W}_{1}^{(t)}(\bm{y},\bm{X}\bm{U},\bm{W}_{1}^{(t-1)},\mathcal{W}_{\setminus 1}^{(t-1)})$ (20)
$=\bm{W}_{1}^{(t-1)}-\eta_{t}\sum_{i\in I_{t}}\partial_{2}\mathcal{L}\left(y_{i},f_{\mathcal{W}^{(t-1)}}\left(\bm{U}^{\top}\bm{x}_{i}\right)\right)\cdot\bm{U}^{\top}\bm{x}_{i}h\left((\bm{W}_{1}^{(t-1)})^{\top}\bm{U}^{\top}\bm{x}_{i};\mathcal{W}_{\setminus 1}^{(t-1)}\right)^{\top}$ (21)
$=\bm{U}^{\top}\bm{W}_{1}^{(t)}(\bm{y},\bm{X},\bm{U}\bm{W}_{1}^{(t-1)},\mathcal{W}_{\setminus 1}^{(t-1)}),$ (22)
where we use $f_{\mathcal{W}^{(t-1)}}\left(\bm{U}^{\top}\bm{x}_{i}\right)=f_{(\bm{U}\bm{W}_{1}^{(t-1)},\mathcal{W}_{\setminus 1}^{(t-1)})}\left(\bm{x}_{i}\right)$ in the last equation.
By Assumption 1 (i) and $\bm{U}\bm{B}=\bm{B}$, it follows that, for each $i\in[m]$,
$y_{i}=g(\bm{B}^{\top}\bm{x}_{i},\varepsilon_{i})=g(\bm{B}^{\top}\bm{U}^{\top}\bm{x}_{i},\varepsilon_{i}).$ (23)
Hence, $(\bm{y},\bm{X})\overset{\rm d}{=}(\bm{y},\bm{X}\bm{U})$ by Assumption 1 (ii). Together with Assumption 4, we obtain
$(\bm{y},\bm{X}\bm{U},\bm{W}_{1}^{(0)},\mathcal{W}_{\setminus 1}^{(0)})\overset{\rm d}{=}(\bm{y},\bm{X},\bm{U}\bm{W}_{1}^{(0)},\mathcal{W}_{\setminus 1}^{(0)}).$ (24)
This implies
$\bm{W}_{1}^{(1)}(\bm{y},\bm{X}\bm{U},\bm{W}_{1}^{(0)},\mathcal{W}_{\setminus 1}^{(0)})\overset{\rm d}{=}\bm{W}_{1}^{(1)}(\bm{y},\bm{X},\bm{U}\bm{W}_{1}^{(0)},\mathcal{W}_{\setminus 1}^{(0)}),$ (25)
which follows from the fact that the identical measurable function of the random elements following the same law has again the same law. As a result, (21) for $t=1$, (25), and the left-orthogonal invariance of $\bm{W}_{1}^{(0)}$ in Assumption 4 give
$\bm{U}\bm{W}_{1}^{(1)}(\bm{y},\bm{X},\bm{W}_{1}^{(0)},\mathcal{W}_{\setminus 1}^{(0)})\overset{\rm d}{=}\bm{W}_{1}^{(1)}(\bm{y},\bm{X},\bm{W}_{1}^{(0)},\mathcal{W}_{\setminus 1}^{(0)}).$ (26)
Also for $t=2$, applying (21) yields
$\bm{U}\bm{W}_{1}^{(2)}\left(\bm{y},\bm{X}\bm{U},\bm{W}_{1}^{(1)}(\bm{y},\bm{X}\bm{U},\bm{W}_{1}^{(0)},\mathcal{W}_{\setminus 1}^{(0)}),\mathcal{W}_{\setminus 1}^{(1)}\right)$ (27)
$=\bm{W}_{1}^{(2)}\left(\bm{y},\bm{X},\bm{U}\bm{W}_{1}^{(1)}(\bm{y},\bm{X}\bm{U},\bm{W}_{1}^{(0)},\mathcal{W}_{\setminus 1}^{(0)}),\mathcal{W}_{\setminus 1}^{(1)}\right)$ (28)
$=\bm{W}_{1}^{(2)}\left(\bm{y},\bm{X},\bm{W}_{1}^{(1)}(\bm{y},\bm{X},\bm{U}\bm{W}_{1}^{(0)},\mathcal{W}_{\setminus 1}^{(0)}),\mathcal{W}_{\setminus 1}^{(1)}\right).$ (29)
Therefore, we have, by (24),
$\bm{U}\bm{W}_{1}^{(2)}(\bm{y},\bm{X},\bm{W}_{1}^{(1)},\mathcal{W}_{\setminus 1}^{(1)})\overset{\rm d}{=}\bm{W}_{1}^{(2)}(\bm{y},\bm{X},\bm{W}_{1}^{(1)},\mathcal{W}_{\setminus 1}^{(1)}),$ (30)
in the same manner as (26). We can immediately generalize this argument to any $t\in\mathbb{N}$ by the recursive argument. This implies the desired property (17). ∎
A.2. Proof of Theorem 1
For the proof of Theorem 1, we prepare a lemma.
Lemma A.1.
Assume the Assumption 1. Let $\bm{z}\in\mathbb{R}^{n}$ be a standard Gaussian random vector. For any $t>0$, there exists a universal constant $c>0$ such that
$\mathbb{P}\left(\left\lvert\sqrt{\frac{\bm{z}^{\top}\bm{P}_{\bm{B}}^{\perp}\bm{z}}{n}}-\sqrt{\frac{n-q^{*}}{n}}\right\rvert>t\right)\leq 2\exp\left(-c\left(nt^{2}\wedge\sqrt{n(n-q^{*})}t\right)\right).$ (31)
Proof of Lemma A.1.
Since $\mathsf{rank}(\bm{B})=q^{*}$ by Assumption 1 (i), we have $\|\bm{P}_{\bm{B}}^{\perp}\|_{F}^{2}=n-q^{*}$ and $\|\bm{P}_{\bm{B}}^{\perp}\|_{\rm op}=1$. Thus, Hanson–Wright inequality [vershynin2018high] implies that, for any $t>0$,
$\mathbb{P}\left(\left\lvert\bm{z}^{\top}\bm{P}_{\bm{B}}\bm{z}-(n-q^{*})\right\rvert>t\right)\leq 2\exp\left(-c\min\left(\frac{t^{2}}{n-q^{*}}\wedge t\right)\right),$ (32)
with some constant $c>0$. Hence, it follows that, for any $t>0$,
$\mathbb{P}\left(\left\lvert\frac{\bm{z}^{\top}\bm{P}_{\bm{B}}\bm{z}}{n-q^{*}}-1\right\rvert>t\right)\leq 2\exp\left(-c\,(n-q^{*})(t^{2}\wedge t)\right),$ (33)
with some constant $c>0$. For $a>0$, we have $|a^{2}-1|=|a+1|\cdot|a-1|\geq|a-1|$ since $a+1>1$. Using this, we obtain, for any $t>0$,
$\mathbb{P}\left(\left\lvert\sqrt{\frac{\bm{z}^{\top}\bm{P}_{\bm{B}}\bm{z}}{n-q^{*}}}-1\right\rvert>t\right)\leq\mathbb{P}\left(\left\lvert\frac{\bm{z}^{\top}\bm{P}_{\bm{B}}\bm{z}}{n-q^{*}}-1\right\rvert>t\right)\leq 2\exp\left(-c\,(n-q^{*})(t^{2}\wedge t)\right),$ (34)
with some constant $c>0$. Change-of-variable from $t$ to $t\sqrt{n/(n-q^{*})}$ completes the proof. ∎
Proof of Theorem 1.
From Proposition 1, we obtain
$\frac{\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}\overset{\rm d}{=}\frac{\bm{P}_{\bm{B}}^{\perp}\bm{z}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{z}\|},$ (35)
where $\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$. Here, $j$-th element of $\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}$ and $\bm{P}_{\bm{B}}^{\perp}\bm{z}$ are given by
$\xi_{j}^{(t)}-\bm{b}_{j}^{\top}(\bm{B}^{\top}\bm{B})^{-1}\bm{B}^{\top}\bm{\xi}^{(t)}\quad\mathrm{and}\quad z_{j}-\bm{b}_{j}^{\top}(\bm{B}^{\top}\bm{B})^{-1}\bm{B}^{\top}\bm{z},$ (36)
respectively. Note that $\bm{b}_{j}=\bm{0}$ under the null by Theorem 3. Thus, for $j\in S^{\mathsf{c}}$, we have
$\sqrt{n}\xi_{j}^{(t)}/\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|\overset{\rm d}{=}\sqrt{n}z_{j}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{z}\|}.$ (37)
Here, since $\mathrm{rank}(\bm{B})=q^{*}$ by Assumption 1 (i), Lemma A.1 implies that $\|\bm{P}_{\bm{B}}^{\perp}\bm{z}\|/\sqrt{n}\overset{\mathrm{p}}{\to}1$ as $n\to\infty$ while $q^{*}=o(n)$. This completes the proof of (7).
Next, we show the uniform convergence. Denote $\sigma_{n}=\sqrt{n}/\|\bm{P}_{\bm{B}}^{\perp}\bm{z}\|$ for convenience. Fix an arbitrary $\epsilon>0$ and define $E_{n}=\left\{|\sigma_{n}-1|<\epsilon\right\}$. From (37), we have
$\Delta_{n}\equiv\sup_{j\in S^{\mathsf{c}}}\ \sup_{u\in\mathbb{R}}\left\lvert\mathbb{P}\left(\frac{\sqrt{n}\,\xi_{j}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}\leq u\right)-\Phi(u)\right\rvert$ $=\sup_{j\in S^{\mathsf{c}}}\ \sup_{u\in\mathbb{R}}\left\lvert\mathbb{P}\left(z_{j}\leq u\sigma_{n}\right)-\Phi(u)\right\rvert$. (38)
Here, we have
$\left\lvert\mathbb{P}\left(z_{j}\leq u\sigma_{n}\right)-\Phi(u)\right\rvert\leq\mathbb{P}(E_{n}^{\mathsf{c}})+\left\lvert\mathbb{P}(\sigma_{n}z_{j}<u\mid E_{n})-\Phi(t)\right\rvert$. (39)
Also, since $\sigma_{n}\in[1-\epsilon,1+\epsilon]$ under $E_{n}$, it follows that
$\mathbb{P}\left(z_{j}<\frac{u}{1+\epsilon}\right)\leq\mathbb{P}\left(\sigma_{n}z_{j}\mid E_{n}\right)\leq\mathbb{P}\left(z_{j}<\frac{u}{1-\epsilon}\right).$ (40)
From this, we have
$\left\lvert\mathbb{P}(\sigma_{n}z_{j}<u\mid E_{n})-\Phi(u)\right\rvert\leq\max\left\{\Phi\left(\frac{u}{1-\epsilon}\right)-\Phi(u),\ \Phi(u)-\Phi\left(\frac{u}{1+\epsilon}\right)\right\}.$ (41)
Taking the supremum, the mean-value theorem gives
$\sup_{u\in\mathbb{R}}\left\lvert\mathbb{P}(\sigma_{n}z_{j}<u\mid E_{n})-\Phi(u)\right\rvert\leq\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{\sqrt{2\pi e}},$ (42)
where we use the fact $\sup_{u\in\mathbb{R}}|u|\phi(u)=1/\sqrt{2\pi e}$. Since this upper bound does not depend on $j\in[n]$, (39) and (42) yield
$\Delta_{n}\leq\mathbb{P}\left(|\sigma_{n}-1|>\epsilon\right)+\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{\sqrt{2\pi e}}.$ (43)
The first term on the right-hand side converges to zero as $n\to\infty$ while $q^{*}=o(n)$ by Lemma A.1, and the second term goes to zero as $\epsilon\downarrow 0$. ∎
A.3. Proof of Thorem 2
Our argument is inspired by the proof of Proposition 3.2 of dai2023scale, but proceeds under weaker conditions on signal strength and dimensionality. While auxiliary lemmas overlap with dai2023scale, we provide full proofs to ensure a self-contained presentation. We prepare some notation for the proof.
  • $I_{u}(v)\equiv\inf\{w\geq 0:\psi(v,w)>u\}$ for any $u>0$ and $v\geq 0$ with the convention $\inf\emptyset=+\infty$.
  • $\tilde{M}_{j}=\mathsf{sign}(z_{j1}z_{j2})\psi(|z_{j1}|,|z_{j2}|)$ where $(z_{11},\ldots,z_{n1},z_{12},\ldots,z_{n2})^{\top}\sim\mathcal{N}(\bm{0},\bm{I}_{2n})$.
  • $\varsigma_{j1}^{(t)}=\sqrt{n}\xi_{j1}^{(t)}/\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}_{1}^{(t)}\|$ and $\varsigma_{j2}^{(t)}=\sqrt{n}\xi_{j2}^{(t)}/\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}_{1}^{(t)}\|$ for any $t\in\mathbb{N}$ and $j\in[n]$.
  • $\breve{M}_{j}=\mathsf{sign}(\varsigma_{j1}^{(t)}\varsigma_{j2}^{(t)})\psi(|\varsigma_{j1}^{(t)}|,|\varsigma_{j2}^{(t)}|)$.
  • $V^{+}(u)=\#\{j\in S^{\mathsf{c}}:\breve{M}_{j}>u\}$ and $V^{-}(u)=\#\{j\in S^{\mathsf{c}}:\breve{M}_{j}<-u\}$.
  • $\tilde{F}(u)=\mathbb{P}(\tilde{M}_{1}>u)$.
  • $\tau_{\alpha}^{\sigma_{n}}$ is the cutoff when we apply Algorithm 1 with $\breve{M}_{j}$ instead of $M_{j}$.
Recall that we defined the original importance statistics as ${M}_{j}=\mathsf{sign}(\xi_{j1}^{(t)}\xi_{j2}^{(t)})\psi(|\xi_{j1}^{(t)}|,|\xi_{j2}^{(t)}|)$.
Lemma A.2 (Zero-symmetry of $\tilde{M}_{j}$).
For each $j\in[n]$ and any Borel measurable function $\psi:[0,\infty)^{2}\to[0,\infty)$, we have
$\tilde{M}_{j}\;\overset{\mathrm{d}}{=}\;-\,\tilde{M}_{j}.$
Proof.
Write $f:\mathbb{R}^{2}\to\mathbb{R}$ for the measurable map
$f(z_{1},z_{2})\;=\;\operatorname{sign}(z_{1}z_{2})\,\psi\bigl(|z_{1}|,|z_{2}|\bigr).$
We observe the oddness under a single-coordinate reflection: for all $z_{1},z_{2}\in\mathbb{R}$,
$f(-z_{1},z_{2})\;=\;-\;f(z_{1},z_{2}),$ (44)
since $\mathsf{sign}((-z_{1})z_{2})=-\mathsf{sign}(z_{1}z_{2})$ while the arguments of $\psi$ are unchanged by taking absolute values.
Since $(z_{j1},z_{j2})^{\top}\sim\mathcal{N}(\bm{0},\bm{I}_{2})$, its law is invariant under the orthogonal reflection
$R:=\begin{pmatrix}-1&0\\ 0&1\end{pmatrix}$,
i.e., $(z_{j1},z_{j2})^{\top}\overset{\mathrm{d}}{=}R(z_{j1},z_{j2})^{\top}=(-z_{j1},z_{j2})^{\top}$. Therefore, for any Borel set $A\subset\mathbb{R}$,
$\mathbb{P}\bigl(\tilde{M}_{j}\in A\bigr)$ $=\mathbb{P}\bigl(f(z_{j1},z_{j2})\in A\bigr)=\mathbb{P}\bigl(f(-z_{j1},z_{j2})\in A\bigr)=\mathbb{P}\bigl(-f(z_{j1},z_{j2})\in A\bigr)$
$=\mathbb{P}\bigl(f(z_{j1},z_{j2})\in-A\bigr)=\mathbb{P}\bigl(\tilde{M}_{j}\in-A\bigr).$
Since this holds for every Borel $A$, the laws of $\tilde{M}_{j}$ and $\tilde{M}_{j}$ coincide; that is, $\tilde{M}_{j}\overset{\mathrm{d}}{=}-\tilde{M}_{j}$. ∎
Lemma A.3 (Invariance under common scaling).
Suppose Assumption 5 holds. For any $C\in\mathbb{R}\setminus\{0\}$, define scaled importance scores for each $j\in[n]$,
$\tilde{\xi}_{j\cdot}^{(t)}=C\,\xi_{j\cdot}^{(t)},\quad M_{j}^{C}=\mathsf{sign}(\tilde{\xi}_{j1}^{(t)}\tilde{\xi}_{j2}^{(t)})\,\psi\bigl(|\tilde{\xi}_{j1}^{(t)}|,|\tilde{\xi}_{j2}^{(t)}|\bigr).$
Let the corresponding cutoff and the set of selected indices be $\tau_{\alpha}^{C},{\hat{S}}_{\alpha}^{C}$. Then
$\tau_{\alpha}^{C}=|C|^{r}\,\tau_{\alpha},\qquad{\hat{S}}_{\alpha}^{C}=\hat{S}_{\alpha}.$
Proof.
At first, the sign factor is invariant since, for each $j\in[n]$,
$\mathsf{sign}\bigl(\tilde{\xi}_{j1}^{(t)}\tilde{\xi}_{j2}^{(t)}\bigr)=\mathsf{sign}\bigl((C\xi_{j1}^{(t)})(C\xi_{j2}^{(t)})\bigr)=\mathsf{sign}(C^{2})\,\mathsf{sign}\bigl(\xi_{j1}^{(t)}\xi_{j2}^{(t)}\bigr)=\mathsf{sign}\bigl(\xi_{j1}^{(t)}\xi_{j2}^{(t)}\bigr).$
Also, since $|\tilde{\xi}_{jk}^{(t)}|=|C|\,|\xi_{jk}^{(t)}|$, Assumption 5 of homogeneity implies that there exists $r>0$ such that
$\psi(|\tilde{\xi}_{j1}^{(t)}|,|\tilde{\xi}_{j2}^{(t)}|)=\psi(|C|\,|\xi_{j1}^{(t)}|,\,|C|\,|\xi_{j2}^{(t)}|)=|C|^{r}\,\psi(|\xi_{j1}^{(t)}|,|\xi_{j2}^{(t)}|).$
Hence, for all $j\in[n]$,
$M_{j}^{C}=|C|^{r}\,M_{j}.$ (45)
Then, we have, for any $u>0$,
$\{j:M_{j}^{C}>u\}=\{j:|C|^{r}M_{j}>u\}=\{j:M_{j}>u/|C|^{r}\},$
$\{j:M_{j}^{C}<-u\}=\{j:|C|^{r}M_{j}<-u\}=\{j:M_{j}<-u/|C|^{r}\}.$
The counterpart $\widehat{\mathsf{FDP}}^{C}$ of $\widehat{\mathsf{FDP}}$ satisfies
$\widehat{\mathsf{FDP}}^{C}(u)=\frac{\#\{j:M_{j}^{C}<-u\}}{\#\{j:M_{j}^{C}>u\}\vee 1}=\frac{\#\{j:M_{j}<-u/|C|^{r}\}}{\#\{j:M_{j}>u/|C|^{r}\}\vee 1}=\widehat{\mathsf{FDP}}\left(\frac{u}{|C|^{r}}\right).$ (46)
From (46), it follows that
$\{u>0:{\widehat{\mathsf{FDP}}}^{C}(u)\leq\alpha\}$ $=\{u>0:\widehat{\mathsf{FDP}}(u/|C|^{r})\leq\alpha\}$
$=\{|C|^{r}v:v>0,\ \widehat{\mathsf{FDP}}(v)\leq\alpha\}.$
Therefore, we obtain
$\tau_{\alpha}^{C}=|C|^{r}\,\tau_{\alpha}.$ (47)
By (45) and (47),
${\hat{S}}_{\alpha}^{C}=\{j:M_{j}^{C}>\tau_{\alpha}^{C}\}=\{j:|C|^{r}M_{j}>|C|^{r}\tau_{\alpha}\}=\{j:M_{j}>\tau_{\alpha}\}=\hat{S}_{\alpha}.$
Corollary A.1.
Under Assumption 5, $(M_{j})_{j\in[n]}$ and $(\breve{M}_{j})_{j\in[n]}$ yield the same selection result after applying Algorithm 1.
Proof of Corollary A.1.
Applying Lemma A.3 with $C=\sqrt{n}/\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}_{1}^{(t)}\|$ proves the claim if $n\geq 1$, $\bm{B}\neq\bm{0}_{n\times q^{*}}$, and $\bm{\xi}_{1}^{(t)}\neq\bm{0}_{n}$. Note that the convergence of $\bm{\xi}_{1}^{(t)}$ to zero is allowed. ∎
Lemma A.4.
Under the assumptions of Theorem 2, as $n\to\infty$ while $q^{*}=o(n)$, we have
$\sup_{u\in\mathbb{R},\,j\in S^{\mathsf{c}}}\left\lvert\mathbb{P}(\breve{M}_{j}>u)-\mathbb{P}(\tilde{M}_{j}>u)\right\rvert\to 0.$ (48)
Proof of Lemma A.4.
Define
$\Delta_{j}=\sup_{u\in\mathbb{R}}|\mathbb{P}(\varsigma_{j1}^{(t)}>u)-\mathbb{P}(z_{j1}>u)|\ \vee\ \sup_{u\in\mathbb{R}}|\mathbb{P}(\varsigma_{j2}^{(t)}>u)-\mathbb{P}(z_{j1}>u)|.$ (49)
Without loss of generality, we assume $u>0$. Thus, by the non-negativeness of $\psi(\cdot,\cdot)$, we have
$\{\breve{M}_{j}>u\}\iff\left(\left\{\psi(|\varsigma_{j1}^{(t)}|,|\varsigma_{j2}^{(t)}|)>u\right\}\cap\{\varsigma_{j1}^{(t)}>0\}\right)\cup\left(\left\{-\psi(|\varsigma_{j1}^{(t)}|,|\varsigma_{j2}^{(t)}|)>u\right\}\cap\{\varsigma_{j1}^{(t)}\leq 0\}\right).$ (50)
Using this and the monotonicity of $\psi(\cdot,\cdot)$, for any $t\in\mathbb{N}$, we have
$\mathbb{P}(\breve{M}_{j}>u)$ $=\mathbb{P}\left(\varsigma_{j2}^{(t)}>I_{u}(\varsigma_{j1}^{(t)}),\,\varsigma_{j1}^{(t)}>0\right)+\mathbb{P}\left(\varsigma_{j2}^{(t)}<-I_{u}(\varsigma_{j1}^{(t)}),\,\varsigma_{j1}^{(t)}<0\right)$ (51)
$\leq\mathbb{P}\left(z_{j2}>I_{u}(\varsigma_{j1}^{(t)}),\,\varsigma_{j1}^{(t)}>0\right)+\mathbb{P}\left(z_{j2}<-I_{u}(\varsigma_{j1}^{(t)}),\,\varsigma_{j1}^{(t)}<0\right)+2\Delta_{j}$ (52)
$=\mathbb{P}\left(\mathsf{sign}(\varsigma_{j1}^{(t)}z_{j2})\psi(|\varsigma_{j1}^{(t)}|,|z_{j2}|)>u\right)+2\Delta_{j}$ (53)
$=\mathbb{P}\left(\varsigma_{j1}^{(t)}>I_{u}(z_{j2}),\,z_{j2}>0\right)+\mathbb{P}\left(\varsigma_{j1}^{(t)}<-I_{u}(z_{j2}),\,z_{j2}<0\right)+2\Delta_{j}$ (54)
$\leq\mathbb{P}\left(z_{j1}>I_{u}(z_{j2}),\,z_{j2}>0\right)+\mathbb{P}\left(z_{j1}<-I_{u}(z_{j2}),\,z_{j2}<0\right)+4\Delta_{j}$ (55)
$=\mathbb{P}(\tilde{M}_{j}>u)+4\Delta_{j},$ (56)
where the first inequality follows from $\varsigma_{j1}^{(t)}\overset{\mathrm{d}}{=}\varsigma_{j2}^{(t)}$ by Assumption 6, and the third equality follows from the symmetry of $\psi(\cdot,\cdot)$. Hence, Theorem 1 implies that
$\sup_{u\in\mathbb{R},\,j\in S^{\mathsf{c}}}\left\lvert\mathbb{P}(\breve{M}_{j}>u)-\mathbb{P}(\tilde{M}_{j}>u)\right\rvert\leq 4\sup_{u\in\mathbb{R},\,j\in S^{\mathsf{c}}}|\Delta_{j}|\to 0,$ (57)
as $n\to\infty$ with $q^{*}=o(n)$. ∎
Lemma A.5.
Let $n_{0}$ be the number of null features. Under the assumptions of Theorem 2, as $n\to\infty$ while $q^{*}=o(n)$, we have
$\sup_{u\in\mathbb{R}}\mathrm{Var}\left(\frac{1}{n_{0}}\sum_{j\in S^{\mathsf{c}}}\mathbbm{1}(\breve{M}_{j}>u)\right)\leq\frac{1}{4n_{0}}+o(1).$ (58)
Proof of Lemma A.5.
We assume $u>0$ without loss of generality. It follows that
$\sup_{u\in\mathbb{R}}\mathrm{Var}\left(\frac{1}{n_{0}}\sum_{j\in S^{\mathsf{c}}}\mathbbm{1}(\breve{M}_{j}>u)\right)$ (59)
$\leq\frac{1}{n_{0}^{2}}\sum_{j\in S^{\mathsf{c}}}\sup_{u\in\mathbb{R}}\mathrm{Var}\left(\mathbbm{1}(\breve{M}_{j}>u)\right)+\frac{1}{n_{0}^{2}}\sum_{j\neq j^{\prime}\in S^{\mathsf{c}}}\sup_{u\in\mathbb{R}}\mathrm{Cov}\left(\mathbbm{1}(\breve{M}_{j}>u),\mathbbm{1}(\breve{M}_{j^{\prime}}>u)\right),$ (60)
where $\mathbbm{1}(\breve{M}_{j}>u)$ is a Bernoulli variable, and its variance is bounded above by $1/4$. Thus, the first term on the right-hand side is upper bounded by $1/(4n_{0})$. For the first term, we have
$\mathrm{Cov}\left(\mathbbm{1}(\breve{M}_{j}>u),\mathbbm{1}(\breve{M}_{j^{\prime}}>u)\right)$ (61)
$=\mathbb{P}(\breve{M}_{j}>u,\breve{M}_{j^{\prime}}>u)-\mathbb{P}(\breve{M}_{j}>u)\mathbb{P}(\breve{M}_{j^{\prime}}>u)$ (62)
$\leq\left\lvert\mathbb{P}(\breve{M}_{j}>u,\breve{M}_{j^{\prime}}>u)-\mathbb{P}(\tilde{M}_{j}>u)^{2}\right\rvert+\left\lvert\mathbb{P}(\breve{M}_{j}>u)\mathbb{P}(\breve{M}_{j^{\prime}}>u)-\mathbb{P}(\tilde{M}_{j}>u)^{2}\right\rvert$, (63)
where the second term on the right-hand side converges to zero uniformly on $j\in S^{\mathsf{c}}$ and $u\in\mathbb{R}$ by Lemma A.4.
Repeating the argument in the proof of Lemma A.4, it follows that
$\mathbb{P}(\breve{M}_{j}>u,\breve{M}_{j^{\prime}}>u)$ $=\mathbb{P}\left(\varsigma_{j^{\prime}2}^{(t)}>I_{u}(\varsigma_{j^{\prime}1}^{(t)}),\,\varsigma_{j2}^{(t)}>I_{u}(\varsigma_{j1}^{(t)}),\,\varsigma_{j^{\prime}1}^{(t)}>0,\,\varsigma_{j1}^{(t)}>0\right)$ (64)
$+\mathbb{P}\left(\varsigma_{j^{\prime}2}^{(t)}>I_{u}(\varsigma_{j^{\prime}1}^{(t)}),\,\varsigma_{j2}^{(t)}<-I_{u}(\varsigma_{j1}^{(t)}),\,\varsigma_{j^{\prime}1}^{(t)}>0,\,\varsigma_{j1}^{(t)}<0\right)$ (65)
$+\mathbb{P}\left(\varsigma_{j^{\prime}2}^{(t)}<-I_{u}(\varsigma_{j^{\prime}1}^{(t)}),\,\varsigma_{j2}^{(t)}>I_{u}(\varsigma_{j1}^{(t)}),\,\varsigma_{j^{\prime}1}^{(t)}<0,\,\varsigma_{j1}^{(t)}>0\right)$ (66)
$+\mathbb{P}\left(\varsigma_{j^{\prime}2}^{(t)}<-I_{u}(\varsigma_{j^{\prime}1}^{(t)}),\,\varsigma_{j2}^{(t)}<-I_{u}(\varsigma_{j1}^{(t)}),\,\varsigma_{j^{\prime}1}^{(t)}<0,\,\varsigma_{j1}^{(t)}<0\right)$ (67)
$\equiv I_{1}+I_{2}+I_{3}+I_{4}.$ (68)
Define
$\Delta=\sup_{j\in S^{\mathsf{c}},\,u\in\mathbb{R}}|\mathbb{P}(\varsigma_{j1}^{(t)}>u)-\mathbb{P}(z_{j1}>u)|\ \vee\ \sup_{j\in S^{\mathsf{c}},\,u\in\mathbb{R}}|\mathbb{P}(\varsigma_{j2}^{(t)}>u)-\mathbb{P}(z_{j1}>u)|.$ (69)
Let $Q(u)=1-\Phi(u)$. For $I_{1}$, we have the following upper bound,
$I_{1}$ $=\mathbb{E}\left[\mathbb{P}(\varsigma_{j^{\prime}2}^{(t)}>I_{u}(x),\,\varsigma_{j2}^{(t)}>I_{u}(y))\mid\varsigma_{j^{\prime}1}^{(t)}=x>0,\,\varsigma_{j1}^{(t)}=y>0\right]$ (70)
$\leq\mathbb{E}\left[\mathbb{P}(z_{j^{\prime}2}>I_{u}(x),\,z_{j2}>I_{u}(y))\mid\varsigma_{j^{\prime}1}^{(t)}=x>0,\,\varsigma_{j1}^{(t)}=y>0\right]+2\Delta$ (71)
$=\mathbb{E}\left[Q(I_{u}(x))Q(I_{u}(y))\mid\varsigma_{j^{\prime}1}^{(t)}=x>0,\,\varsigma_{j1}^{(t)}=y>0\right]+2\Delta.$ (72)
Similarly, we can upper bound $I_{2}$, $I_{3}$, and $I_{4}$. Combining the four upper bounds together, we obtain an upper bound on $\mathbb{P}(\breve{M}_{j}>u,\breve{M}_{j^{\prime}}>u)$ as
$\mathbb{P}\left(\mathsf{sign}(z_{j2}\varsigma_{j1}^{(t)})\psi(|z_{j2}|,|\varsigma_{j1}^{(t)}|)>u,\ \mathsf{sign}(z_{j^{\prime}2}\varsigma_{j^{\prime}1}^{(t)})\psi(|z_{j^{\prime}2}|,|\varsigma_{j^{\prime}1}^{(t)}|)>u\right)+8\Delta.$ (73)
We can further decompose this into four terms as (66) by conditioning the signs of $z_{j2}$ and $z_{j^{\prime}2}$, and repeat the upper bound (71). This leads to
$\mathbb{P}(\breve{M}_{j}>u,\breve{M}_{j^{\prime}}>u)\leq\mathbb{P}(\tilde{M}_{j}>u)^{2}+16\Delta.$ (74)
Similarly, we can show the corresponding lower bound. Since $\Delta\to 0$ by Lemma A.4, the covariance in (61) converges to zero. ∎
Lemma A.6.
Under the assumptions of Theorem 2, we have, as $n\to\infty$ while $q^{*}=o(n)$,
$\sup_{u\in\mathbb{R}}\left\lvert\frac{1}{n_{0}}\sum_{j\in S^{\mathsf{c}}}\mathbbm{1}(\breve{M}_{j}>u)-\mathbb{P}(\tilde{M}_{1}>u)\right\rvert\overset{\mathrm{p}}{\to}0.$ (75)
Proof of Lemma A.6.
We have
$\sup_{u\in\mathbb{R}}\left\lvert\frac{1}{n_{0}}\sum_{j\in S^{\mathsf{c}}}\mathbbm{1}(\breve{M}_{j}>u)-\mathbb{P}(\tilde{M}_{1}>u)\right\rvert$ (76)
$\leq\sup_{u\in\mathbb{R}}\left\lvert\frac{1}{n_{0}}\sum_{j\in S^{\mathsf{c}}}\left\{\mathbbm{1}(\breve{M}_{j}>u)-\mathbb{P}(\breve{M}_{j}>u)\right\}\right\rvert+\sup_{u\in\mathbb{R}}\left\lvert\frac{1}{n_{0}}\sum_{j\in S^{\mathsf{c}}}\mathbb{P}(\breve{M}_{j}>u)-\mathbb{P}(\tilde{M}_{1}>u)\right\rvert$, (77)
where the second term on the right-hand side converges to zero by Lemma A.4. Also, Chebyshev’s inequality yields, for any $v\in\mathbb{R}$,
$\sup_{u\in\mathbb{R}}\,\mathbb{P}\left(\left\lvert\frac{1}{n_{0}}\sum_{j\in S^{\mathsf{c}}}\left\{\mathbbm{1}(\breve{M}_{j}>u)-\mathbb{P}(\breve{M}_{j}>u)\right\}\right\rvert>v\right)\leq\frac{1}{v^{2}}\sup_{u\in\mathbb{R}}\mathrm{Var}\left(\frac{1}{n_{0}}\sum_{j\in S^{\mathsf{c}}}\mathbbm{1}(\breve{M}_{j}>u)\right),$ (78)
where the supremum of variance converges to zero by Lemma A.5. This completes the proof. ∎
Lemma A.7.
Under the assumptions of Theorem 2, there exists a constant $\delta_{0}=\delta_{0}(\alpha,c,\theta,\rho,\pi_{0})>0$ such that
$\mathbb{P}\bigl(\,\tilde{F}(\tau_{\alpha}^{\sigma_{n}})\geq\delta_{0}\,\bigr)\;\longrightarrow\;1.$
Moreover, one can take explicitly
$\delta_{\ast}:=\frac{(1-\theta)c}{2\pi_{0}},\qquad U:=\frac{(\alpha\rho-(1-\rho))\,\theta c}{\pi_{0}(1-\alpha)},\qquad\delta_{0}:=\frac{\delta_{\ast}+U}{2}\in(0,1).$
Proof.
Define the generalized inverse $\tilde{F}^{\leftarrow}(\delta):=\inf\{u\geq 0:\ \tilde{F}(u)<\delta\}$ (so $\tilde{F}(\tilde{F}^{\leftarrow}(\delta))\geq\delta$). Note that $\tau_{\alpha}^{\sigma_{n}}$, $V^{+}$, and $V^{-}$ are defined at the top of this section.
Step 1: An upper bound for $\tilde{F}(u_{K_{n}})$. By Assumption 7, $S^{\pm}(u_{K_{n}})\geq\theta K_{n}$ with probability 1o(1)1-o(1); hence the number of nulls among the top $K_{n}$ magnitudes satisfies
$V^{+}(u_{K_{n}})+V^{-}(u_{K_{n}})\;\leq\;(1-\theta)K_{n}.$
By Lemma A.6, $V^{+}(u)+V^{-}(u)=2n_{0}\tilde{F}(u)+o_{p}(n_{0})$ uniformly in $u$. Since $n_{0}$ and $n$ are of the same order,
$2n_{0}\tilde{F}(u_{K_{n}})\;\leq\;(1-\theta)K_{n}+o_{p}(n),$
and dividing both sides by $n$ yields
$\tilde{F}(u_{K_{n}})\;\leq\;\delta_{\ast}+o_{p}(1).$ (79)
Step 2: Construct a subthreshold $u_{0}$. Let $\delta_{0}:=(\delta_{\ast}+U)/2$, which satisfies $\delta_{\ast}<\delta_{0}<U$ by (13). By (79) and the monotonicity of $\tilde{F}$, we have $u_{0}:=\tilde{F}^{\leftarrow}(\delta_{0})<u_{K_{n}}$ with probability 1o(1)1-o(1). Hence, by monotonicity again,
$S^{\pm}(u_{0})\;\geq\;S^{\pm}(u_{K_{n}})\;\geq\;\theta K_{n},$ (80)
with probability approaching one. Moreover, Assumption 7 (applied for all $u\leq u_{K_{n}}$) gives
$S^{+}(u_{0})\;\geq\;\rho\,S^{\pm}(u_{0}),\qquad S^{-}(u_{0})\;\leq\;(1-\rho)\,S^{\pm}(u_{0}),$ (81)
with probability approaching one. Lemma A.6 implies
$V^{+}(u_{0})\;=\;n_{0}\,\delta_{0}+o_{p}(n_{0}).$ (82)
Step 3: Upper bound for $\widehat{\mathsf{FDP}}^{\sigma_{n}}(u_{0})$. Let $\widehat{\mathsf{FDP}}^{\sigma_{n}}(u)$ denote the version of $\widehat{\mathsf{FDP}}(u)$ with $\breve{M}_{j}$ substituted for $M_{j}$ in the definition. Using (80)–(82), Lemma A.2, and $K_{n}=cn+o(n)$,
$\widehat{\mathsf{FDP}}^{\sigma_{n}}(u_{0})=\frac{1+V^{-}(u_{0})+S^{-}(u_{0})}{V^{+}(u_{0})+S^{+}(u_{0})}\;\leq\;\frac{1+n_{0}\delta_{0}+(1-\rho)S^{\pm}(u_{0})+o_{p}(n)}{n_{0}\delta_{0}+\rho S^{\pm}(u_{0})+o_{p}(n)}.$
Divide numerator and denominator by $n$ and pass to $\limsup$ using $n_{0}/n\to\pi_{0}$ and (80):
$\limsup_{n\to\infty}\ \widehat{\mathsf{FDP}}^{\sigma_{n}}(u_{0})\;\leq\;\frac{\pi_{0}\delta_{0}+(1-\rho)\theta c}{\pi_{0}\delta_{0}+\rho\theta c}\;=:\;\Psi(\delta_{0}).$
By the definition of $U$ and the equivalence
$\Psi(\delta)\leq\alpha\ \Longleftrightarrow\ \pi_{0}(1-\alpha)\,\delta\leq(\alpha\rho-(1-\rho))\,\theta c,$
we have $\Psi(\delta_{0})<\alpha$ because $\delta_{0}<U$. Hence there exists $\varepsilon>0$ such that
$\mathbb{P}\bigl(\,\widehat{\mathsf{FDP}}^{\sigma_{n}}(u_{0})\leq\alpha-\varepsilon\,\bigr)\ \to\ 1.$ (83)
Step 4: Compare $\tau_{\alpha}^{\sigma_{n}}$ to $u_{0}$. By definition $\tau_{\alpha}^{\sigma_{n}}:=\inf\{u>0:\widehat{\mathsf{FDP}}^{\sigma_{n}}(u)\leq\alpha\}$. From (83) we obtain $\tau_{\alpha}^{\sigma_{n}}\leq u_{0}$ with probability 1o(1)1-o(1). Since $\tilde{F}$ is nonincreasing and $\tilde{F}(\tilde{F}^{\leftarrow}(\delta_{0}))\geq\delta_{0}$, we conclude
$\tilde{F}(\tau_{\alpha}^{\sigma_{n}})\ \geq\ \tilde{F}(u_{0})\ \geq\ \delta_{0},$
with probability 1o(1)1-o(1), which proves the claim. ∎
Proof of Theorem 2.
To begin with, we have, by Corollary A.1,
$\mathsf{FDR}$ $=\mathbb{E}\left[\frac{\#\{j\in S^{\mathsf{c}}:\breve{M}_{j}>\tau_{\alpha}^{\sigma_{n}}\}}{\#\{j:\breve{M}_{j}>\tau_{\alpha}^{\sigma_{n}}\}\vee 1}\right]=\mathbb{E}\left[\frac{V^{-}(\tau_{\alpha}^{\sigma_{n}})}{V^{+}(\tau_{\alpha}^{\sigma_{n}})+S^{+}(\tau_{\alpha}^{\sigma_{n}})}\right]$ (84)
$\leq\mathbb{E}\left[\frac{1+V^{-}(\tau_{\alpha}^{\sigma_{n}})+S^{-}(\tau_{\alpha}^{\sigma_{n}})}{V^{+}(\tau_{\alpha}^{\sigma_{n}})+S^{+}(\tau_{\alpha}^{\sigma_{n}})}+\frac{\left\lvert V^{+}(\tau_{\alpha}^{\sigma_{n}})-V^{-}(\tau_{\alpha}^{\sigma_{n}})\right\rvert}{V^{+}(\tau_{\alpha}^{\sigma_{n}})+S^{+}(\tau_{\alpha}^{\sigma_{n}})}\right]$ (85)
$\leq\alpha+\mathbb{E}\left[\frac{\left\lvert V^{+}(\tau_{\alpha}^{\sigma_{n}})-V^{-}(\tau_{\alpha}^{\sigma_{n}})\right\rvert}{V^{+}(\tau_{\alpha}^{\sigma_{n}})+S^{+}(\tau_{\alpha}^{\sigma_{n}})}\right],$ (86)
where the last inequality follows from $\widehat{\mathsf{FDP}}^{\sigma_{n}}(\tau_{\alpha}^{\sigma_{n}})\leq\alpha$ by construction. Since $\mathsf{FDR}=0$ when $V^{+}(\tau_{\alpha}^{\sigma_{n}})+S^{+}(\tau_{\alpha}^{\sigma_{n}})=0$, we consider the case $V^{+}(\tau_{\alpha}^{\sigma_{n}})+S^{+}(\tau_{\alpha}^{\sigma_{n}})\geq 1$. By Lemma A.6 and Lemma A.2, we have, as $n\to\infty$,
$\frac{1}{n_{0}}\left\lvert V^{+}(\tau_{\alpha}^{\sigma_{n}})-V^{-}(\tau_{\alpha}^{\sigma_{n}})\right\rvert\leq\frac{1}{n_{0}}\left\lvert V^{+}(\tau_{\alpha}^{\sigma_{n}})-n_{0}\tilde{F}(\tau_{\alpha}^{\sigma_{n}})\right\rvert+\frac{1}{n_{0}}\left\lvert V^{-}(\tau_{\alpha}^{\sigma_{n}})-n_{0}\tilde{F}(\tau_{\alpha}^{\sigma_{n}})\right\rvert=o_{p}(1).$ (87)
Also, since $n_{0}^{-1}V^{+}(\tau_{\alpha}^{\sigma_{n}})=\tilde{F}(\tau_{\alpha}^{\sigma_{n}})+o_{p}(1)$ by Lemma A.6, Lemma A.7 implies that, as $n\to\infty$,
$\frac{1}{n_{0}}V^{+}(\tau_{\alpha}^{\sigma_{n}})\geq\delta_{0}+o_{p}(1).$ (88)
Therefore, from (87) and (88), we have
$\frac{\left\lvert V^{+}(\tau_{\alpha}^{\sigma_{n}})-V^{-}(\tau_{\alpha}^{\sigma_{n}})\right\rvert}{V^{+}(\tau_{\alpha}^{\sigma_{n}})+S^{+}(\tau_{\alpha}^{\sigma_{n}})}=o_{p}(1).$ (89)
Since the left-hand side is bounded by one, (86), (89), and the bounded convergence theorem yield $\mathsf{FDR}\leq\alpha+o(1)$. ∎
A.4. A necessary and sufficient condition for the conditional null
For probability measures $\mu$ and $\nu$ on $\mathbb{R}$, let $W_{1}(\mu,\nu)$ denote the $1$-Wasserstein distance and $d_{\mathrm{TV}}(\mu,\nu)$ the total variation distance. Write $\mathsf{BL}_{1}:=\{\varphi:\mathbb{R}\to\mathbb{R}:\ \|\varphi\|_{\infty}\leq 1,\ \mathrm{Lip}(\varphi)\leq 1\}$ for bounded $1$-Lipschitz functions.
We consider the multi-index model
$y\;=\;g(\bm{B}^{\top}\bm{x},\ \varepsilon),\qquad\bm{x}\in\mathbb{R}^{n},\ y\in\mathbb{R},\ \bm{B}\in\mathbb{R}^{n\times{q^{*}}},$
with $\varepsilon\perp\!\!\!\perp\bm{x}$, $1\leq{q^{*}}<n$, and $\mathrm{rank}(\bm{B})={q^{*}}$. Let the $j$-th row be $\bm{b}_{j}^{\top}\in\mathbb{R}^{q^{*}}$, and set $\bm{u}:=\bm{B}^{\top}\bm{x}\in\mathbb{R}^{q^{*}}$.
Assumption 8 (Minimal thickness of the projected regressor).
There exists a nonempty open set $O\subset\mathbb{R}^{q^{*}}$ such that the law of $\bm{u}$ admits a Lebesgue density strictly positive on $O$.
Assumption 9 (Local kernel Lipschitzness in $W_{1}$).
Let $K(u):=\mathcal{L}(y\mid\bm{u}=u)$ be the conditional law (a stochastic kernel). There exists $L>0$ such that
$W_{1}\big(K(u),K(u^{\prime})\big)\ \leq\ L\,\|u-u^{\prime}\|\qquad\text{for all }u,u^{\prime}\in O.$
Assumption 10 (Bounded-Lipschitz nondegeneracy).
For every nonzero $\bm{v}\in\mathbb{R}^{q^{*}}$ there exist $\varphi\in\mathsf{BL}_{1}$ and a measurable set $A_{\bm{v}}\subset O$ with positive Lebesgue measure such that the directional derivative $D_{\bm{v}}m_{\varphi}(u)$ exists for Lebesgue-a.e. $u\in A_{\bm{v}}$ and is not a.e. zero on $A_{\bm{v}}$, where $m_{\varphi}(u)\equiv\mathbb{E}[\varphi(y)\mid\bm{u}=u]$.
Assumption 11 (Conditional thickness of $x_{j}$ given $\bm{x}_{-j}$).
For the fixed index $j\in\{1,\dots,n\}$ under consideration, write $\bm{x}=(x_{j},\bm{x}_{-j})$. For almost every $\bm{a}_{-j}\in\mathbb{R}^{n-1}$, the conditional law $\mathcal{L}(x_{j}\mid\bm{x}_{-j}=\bm{a}_{-j})$ has a Lebesgue density that is strictly positive on some nonempty open set $K(\bm{a}_{-j})\subset\mathbb{R}$.
Definition 3 (Conditional null for variable $j$).
We say that the conditional null holds for the index $j\in[n]$ if
$y\perp\!\!\!\perp x_{j}\mid\bm{x}_{-j}.$
Lemma A.8.
Under Assumption 9, for every $\varphi\in\mathsf{BL}_{1}$, the map $m_{\varphi}(u)=\mathbb{E}[\varphi(y)\mid\bm{u}=u]$ is $L$-Lipschitz on $O$. In particular, $m_{\varphi}$ is differentiable almost everywhere on $O$.
Proof.
By the Kantorovich–Rubinstein duality for $W_{1}$ on Polish spaces,
$|m_{\varphi}(u)-m_{\varphi}(u^{\prime})|=\Big|\int\varphi\,dK(u)-\int\varphi\,dK(u^{\prime})\Big|\leq W_{1}(K(u),K(u^{\prime}))\leq L\|u-u^{\prime}\|,$
since $\varphi\in\mathsf{BL}_{1}$ is $1$-Lipschitz and bounded. Rademacher’s theorem ensures almost-everywhere differentiability of Lipschitz maps $m_{\varphi}:O\to\mathbb{R}$. ∎
Lemma A.9.
Fix $j\in\{1,\dots,n\}$. Under Assumptions 8, 11, and the conditional null for $j$ in Definition 3, fix $\varphi\in\mathsf{BL}_{1}$ and write
$\bm{w}:=\sum_{k\neq j}\bm{b}_{k}x_{k},\qquad\bm{v}:=\bm{b}_{j}x_{j},$
so that $\bm{u}=\bm{w}+\bm{v}$. Then for almost every $\bm{a}_{-j}\in\mathbb{R}^{n-1}$, there exists a nonempty open set
$G(\bm{a}_{-j})\ \subset\ \bm{w}(\bm{a}_{-j})+\mathrm{span}\{\bm{b}_{j}\}$
such that $m_{\varphi}(u)$ is almost everywhere constant on $G(\bm{a}_{-j})$.
Proof.
By $\bm{x}\perp\!\!\!\perp\varepsilon$ and the definition of $K$,
$\mathbb{E}[\varphi(y)\mid x_{j},\bm{x}_{-j}]\;=\;\mathbb{E}[\varphi(y)\mid\bm{u}]\;=\;m_{\varphi}(\bm{u}).$
The conditional null $y\perp\!\!\!\perp x_{j}\mid\bm{x}_{-j}$ implies
$\mathbb{E}[\varphi(y)\mid x_{j},\bm{x}_{-j}]=\mathbb{E}[\varphi(y)\mid\bm{x}_{-j}],$
hence
$m_{\varphi}(\bm{w}+\bm{v})=h_{\varphi}(\bm{x}_{-j})\quad\text{a.s.}$
for some measurable $h_{\varphi}(\cdot)$.
By Assumption 11, for almost every $\bm{a}_{-j}$ the conditional support of $x_{j}\mid\bm{x}_{-j}=\bm{a}_{-j}$ contains a nonempty open set $K(\bm{a}_{-j})\subset\mathbb{R}$. Consider the linear map
[…]
Its image is the subspace $\mathrm{span}\{\bm{b}_{j}\}$, and $T$ is open onto its image in finite dimensions. Therefore, for almost every $\bm{a}_{-j}$, the image
$H(\bm{a}_{-j}):=T\bigl(K(\bm{a}_{-j})\bigr)$
is a nonempty open set inside $\mathrm{span}\{\bm{b}_{j}\}$ (with respect to the subspace topology). Consequently,
$G(\bm{a}_{-j}):=\bm{w}(\bm{a}_{-j})+H(\bm{a}_{-j})$
is a nonempty open set within the affine subspace $\bm{w}(\bm{a}_{-j})+\mathrm{span}\{\bm{b}_{j}\}$ on which $m_{\varphi}$ is almost everywhere constant (equal to $h_{\varphi}(\bm{a}_{-j})$). ∎
Theorem 3.
Fix $j\in\{1,\dots,n\}$. Under Assumptions 8, 9, 10, and $1$, the following are equivalent:
${\,y\perp\!\!\!\perp x_{j}\mid\bm{x}_{-j}\,}\qquad\Longleftrightarrow\qquad{\,\bm{b}_{j}=\bm{0}\,}.$
Proof.
($\Leftarrow$) If $\bm{b}_{j}=\bm{0}$, then
$\bm{u}=\sum_{k\neq j}\bm{b}_{k}x_{k}$
is $\sigma(\bm{x}_{-j})$-measurable. By $\bm{x}\perp\!\!\!\perp\bm{\varepsilon}$, the conditional law
$\mathcal{L}(y\mid x_{j},\bm{x}_{-j})=\mathcal{L}\bigl(g(\bm{u},\varepsilon)\mid x_{j},\bm{x}_{-j}\bigr)=K(\bm{u})$
does not depend on $x_{j}$, hence $y\perp\!\!\!\perp x_{j}\mid\bm{x}_{-j}$.
($\Rightarrow$) Suppose $y\perp\!\!\!\perp x_{j}\mid\bm{x}_{-j}$. Fix $\varphi\in\mathsf{BL}_{1}$. By Lemma A.9, for almost every $\bm{a}_{-j}$, $m_{\varphi}$ is almost everywhere constant on a nonempty open set inside the affine subspace $\bm{w}(\bm{a}_{-j})+\mathrm{span}\{\bm{b}_{j}\}$. By Assumption 8, these affine pieces intersect $O$ on sets of positive Lebesgue measure in $\mathbb{R}^{q^{*}}$; by Lemma A.8, $m_{\varphi}$ is locally Lipschitz on $O$, hence (Rademacher) directionally differentiable almost everywhere on $O$. Consequently,
$D_{\bm{v}}m_{\varphi}(u)=0\quad\text{for Lebesgue-a.e.\ }u\in O\ \text{ and all }\bm{v}\in\mathrm{span}\{\bm{b}_{j}\}.$
If $\mathrm{span}\{\bm{b}_{j}\}\neq\{\bm{0}\}$, then there exists a nonzero $\bm{v}\in\mathrm{span}\{\bm{b}_{j}\}$. Assumption 10 then yields some $\tilde{\varphi}\in\mathsf{BL}_{1}$ and a measurable $A_{\bm{v}}\subset O$ of positive Lebesgue measure such that $D_{\bm{v}}m_{\tilde{\varphi}}(u)$ exists for Lebesgue-a.e. $u\in A_{\bm{v}}$ and is not almost everywhere zero on $A_{\bm{v}}$, which contradicts the conclusion above (applied to $\tilde{\varphi}$). Thus necessarily $\mathrm{span}\{\bm{b}_{j}\}=\{\bm{0}\}$, i.e., $\bm{b}_{j}=\bm{0}$. ∎
Remark A.1 (TV-variant for classification).
Assumption 9 can be replaced by the TV version $d_{\mathrm{TV}}(K(u),K(u^{\prime}))\leq L\|u-u^{\prime}\|$ on $O$; then $|m_{\varphi}(u)-m_{\varphi}(u^{\prime})|\leq d_{\mathrm{TV}}(K(u),K(u^{\prime}))$ for $\varphi\in\mathsf{BL}_{1}$, and Lemma A.8 and Theorem 3 remain valid with the same proof. In particular, for the ordinal binary classification model $y=\mathbbm{1}{\{h(u)+\varepsilon>0\}}$ (e.g., logistic regression and Probit model) with $q^{*}=1$, $h(\cdot)$ locally Lipschitz and $\varepsilon$ independent with bounded density, the kernel is TV-Lipschitz and Assumption 10 holds with $\varphi(y)=y$ whenever the class-probability $p(u)=\mathbb{P}(\varepsilon>-h(u))$ is not a.e. flat in any nonzero direction.
Appendix B On the right-orthogonal invariance
In this section, we delineate what kinds of random designs fall into the class of right-orthogonally invariant (ROI) matrices, and what kinds do not. ROI is sometimes assumed in the literature of approximate message passing algorithms in the proportional asymptotics where $n$ and $m$ diverge with $m/n\to\delta\in(0,\infty)$ [rangan2019vector, li2023spectrum]. Subsequently, we discuss in what sense the $\bm{B}$-ROI in Assumption 1 (ii) is relaxed.
B.1. Definition and basic consequences
Definition 4 (Right-orthogonal invariance).
A random matrix $\bm{X}\in\mathbb{R}^{m\times n}$ is ROI if
$\bm{X}\overset{\rm d}{=}\bm{X}\bm{U}\qquad(\forall\,\bm{U}\in O(n)).$
If $\mathbb{E}\|\bm{X}\|_{F}^{2}<\infty$, then ROI implies the isotropy of the column Gram:
$\mathbb{E}[\bm{X}^{\top}\bm{X}]=c\,\bm{I}_{n},\qquad c=\frac{1}{n}\,\mathbb{E}\|\bm{X}\|_{F}^{2},$ (90)
which is a necessary (but not sufficient) condition for ROI. There are several closure properties.
Lemma B.1 (Left-multiplicative closure).
Let $\bm{A}$ and $\bm{Z}$ be independent random matrices. If $\bm{Z}$ is ROI, then $\bm{X}:=\bm{A}\bm{Z}$ is ROI.
Lemma B.2 (Right Haar mixer).
For any (possibly deterministic) $\bm{Y}$ and $\bm{Q}\sim\mathrm{Haar}(O(n))$ independent of $\bm{Y}$, $\bm{X}:=\bm{Y}\bm{Q}$ is ROI.
Lemma B.3 (Orthogonally conjugate mixture).
Let $\bm{\Lambda}$ be a symmetric positive-definite random matrix, and suppose $(\bm{X}\bm{U}\mid\bm{\Lambda})\overset{\rm d}{=}(\bm{X}\mid\bm{U}^{\top}\bm{\Lambda}\bm{U})$ and $\bm{\Lambda}\overset{\rm d}{=}\bm{U}^{\top}\bm{\Lambda}\bm{U}$ for all $\bm{U}$. Then the marginal $\bm{X}$ is ROI.
B.2. Canonical ROI examples
Denote the Stiefel manifold $V_{n,r}\equiv\{\bm{W}:\bm{W}^{\top}\bm{W}=\bm{I}_{r}\}$.
(E1) Matrix-normal with isotropic columns. If $\bm{X}\sim\mathcal{MN}(0,\bm{\Sigma}_{\mathrm{row}},\bm{I}_{n})$, then $\bm{X}$ is ROI. Conversely, $\mathcal{MN}(0,\bm{\Sigma}_{\mathrm{row}},\bm{\Sigma}_{\mathrm{col}})$ with $\bm{\Sigma}_{\mathrm{col}}\not\propto\bm{I}_{n}$ is not ROI.
(E2) Elliptical rows (after whitening). If each row is elliptical $\bm{x}_{i}=\bm{\Sigma}^{1/2}\bm{z}_{i}$ with $\bm{z}_{i}$ spherically symmetric, then $\bm{Z}:=\bm{X}\bm{\Sigma}^{-1/2}$ is ROI.
(E3) Spiked with Haar loadings plus isotropic noise. Let $\bm{X}=\alpha\bm{V}\bm{W}^{\top}+\bm{E}$, where $\bm{W}\in V_{n,r}$ is Haar and $\bm{E}$ is ROI (e.g., i.i.d. Gaussian). Then $\bm{X}$ is ROI (by left Haar-invariance of $\bm{W}$ and Lemma B.1).
(E4) Linear multi-layer with an ROI rightmost factor. If $\bm{X}=\bm{X}_{1}\bm{X}_{2}\cdots\bm{X}_{L}$ with $\bm{X}_{L}$ i.i.d. standard Gaussian, then $\bm{X}$ is ROI (Lemma B.1).
(E5) VAR with orthogonally invariant covariance mixing. With $\bm{X}_{i,\cdot}=\sum_{k=1}^{\nu}\alpha_{k}\bm{X}_{i-k,\cdot}+\bm{\epsilon}_{i}$ and $\bm{\epsilon}_{i}\mid\bm{\Sigma}\sim\mathcal{N}(0,\bm{\Sigma})$, $\bm{\Sigma}\sim\mathrm{InvWishart}(\bm{I}_{n})$, one has $(\bm{X}\bm{U}\mid\bm{\Sigma})\overset{\rm d}{=}(\bm{X}\mid\bm{U}^{\top}\bm{\Sigma}\bm{U})$ and $\bm{\Sigma}$ is orthogonally invariant, hence $\bm{X}$ is ROI by Lemma B.3.
(E6) Stiefel-uniform columns and random right projection. If $\bm{Q}\in V_{n,p}$ is uniform and $\bm{X}=s\,\bm{Q}$, then $\bm{X}\bm{U}\overset{\rm d}{=}\bm{X}$. More generally, for any $\bm{Y}$ independent of $\bm{Q}\simHaar$, $\bm{X}=\bm{Y}\bm{Q}$ is ROI (Lemma B.2).
B.3. Non-ROI archetypes (counterexamples)
(N1) Anisotropic Gaussian across columns. $\bm{x}_{i}\sim\mathcal{N}(0,\bm{\Sigma}_{\mathrm{col}})$ with $\bm{\Sigma}_{\mathrm{col}}\not\propto\bm{I}_{n}$ violates (90).
(N2) Columnwise scaling heterogeneity. $\bm{X}=\bm{Z}\bm{D}$ with i.i.d. isotropic $\bm{Z}$ and non-scalar diagonal $\bm{D}$ has $\mathbb{E}[\bm{X}^{\top}\bm{X}]=\mathbb{E}[\bm{Z}^{\top}\bm{Z}]\bm{D}^{2}\not\propto\bm{I}_{n}$.
(N3) Toeplitz/AR(1) column covariance. $\bm{\Sigma}_{\mathrm{col}}=(\rho^{|i-j|})$ breaks ROI already at the second moment.
(N4) Rademacher i.i.d. entries. Invariance holds only for the finite hyperoctahedral group (sign flips and permutations), not for all $\bm{U}\in O(n)$.
(N5) Blockwise variance mixtures across columns. Different column blocks having different scales violate (90).
B.4. Relaxing ROI to the stabilizer
We relax the ROI assumption to invariance under the stabilizer
$\mathcal{G}_{\bm{B}}:=\{\,\bm{U}\in O(n):\,\bm{U}\bm{B}=\bm{B}\,\}.$
Definition 5 ($\bm{B}$-ROI).
A random matrix $\bm{X}\in\mathbb{R}^{m\times n}$ is $\bm{B}$-ROI if $\bm{X}\overset{\rm d}{=}\bm{X}\bm{U}$ for all $\bm{U}\in\mathcal{G}_{\bm{B}}$.
Let $r=\mathsf{rank}(\bm{B})$ and take $\bm{Q}=[\bm{Q}_{B}\ \bm{Q}_{\perp}]$ with $\mathsf{Col}(\bm{Q}_{B})=\mathsf{Col}(\bm{B})$. Then $\mathcal{G}_{\bm{B}}=\{\,\bm{Q}\,\mathrm{diag}(\bm{I}_{r},\bm{R})\,\bm{Q}^{\top}:\ \bm{R}\in O(n-r)\,\}$ and $\bm{U}\bm{P}_{\bm{B}}^{\perp}=\bm{P}_{\bm{B}}^{\perp}\bm{U}$ for all $\bm{U}\in\mathcal{G}_{\bm{B}}$.
How much weaker?
If $\mathbb{E}\|\bm{X}\|_{F}^{2}<\infty$ and we write $\bm{Q}^{\top}\,\mathbb{E}[\bm{X}^{\top}\bm{X}]\,\bm{Q}=\begin{pmatrix}\bm{A}&\bm{C}\\ \bm{C}^{\top}&\bm{D}\end{pmatrix}$, then $\bm{B}$-ROI forces
$\bm{C}=\mathbf{0},\qquad\bm{D}=c\,\bm{I}_{n-r},$ (91)
while $\bm{A}\in\mathbb{R}^{r\times r}$ is arbitrary. In contrast, ROI requires $\mathbb{E}[\bm{X}^{\top}\bm{X}]=c\,\bm{I}_{n}$. Thus $\bm{B}$-ROI is strictly weaker unless $r=0$ (then it coincides with ROI). At the distributional level, $\bm{B}$-ROI is equivalent to: for all $\bm{U}\in O(n-r)$,
$\big(\bm{X}\bm{P}_{\bm{B}},\ \bm{X}\bm{P}_{\bm{B}}^{\perp}\big)\ \overset{\rm d}{=}\ \big(\bm{X}\bm{P}_{\bm{B}},\ \bm{X}\bm{P}_{\bm{B}}^{\perp}\bm{U}\big),$ (92)
i.e. the conditional law of $\bm{X}\bm{P}_{\bm{B}}^{\perp}$ given $\bm{X}\bm{P}_{\bm{B}}$ is ROI.
New important examples under $\bm{B}$-ROI.
  • Fixed-loading spike + isotropic noise: $\bm{X}=\bm{F}\bm{\Lambda}^{\top}+\bm{E}$ with $\mathsf{Col}(\bm{\Lambda})=\mathsf{Col}(\bm{B})$ and $\bm{E}$ isotropic on $\mathsf{Col}(\bm{B})^{\perp}$. Here $\bm{\Lambda}$ may be deterministic (no Haar randomness needed).
  • Anisotropy/discreteness only along $\mathsf{Col}(\bm{B})$: $\bm{X}=\bm{Z}\bm{\Sigma}^{1/2}$ with $\bm{Q}^{\top}\bm{\Sigma}\bm{Q}=\mathrm{diag}(\bm{\Sigma}_{B},\ \sigma^{2}\bm{I}_{n-r})$, where $\bm{\Sigma}_{B}$ is arbitrary SPD; or $\bm{X}\bm{P}_{\bm{B}}$ is discrete/binary while $\bm{X}\bm{P}_{\bm{B}}^{\perp}$ is continuous isotropic (Gaussian/tt/elliptical).
  • Row dependence with conjugate mixing on the complement: VAR-type rows with innovations covariance $\bm{\Sigma}$ satisfying $\bm{Q}^{\top}\bm{\Sigma}\bm{Q}=\mathrm{diag}(\bm{\Sigma}_{B},\ \sigma^{2}\bm{I}_{n-r})$.
  • Partial random right projection: $\bm{X}=\bm{Y}\bm{Q}_{\perp}$ with $\bm{Q}_{\perp}\in V_{n,n-r}$ uniform and $\mathsf{Col}(\bm{Q}_{\perp})=\mathsf{Col}(\bm{B})^{\perp}$.
Remark.
When $r=n$, $\mathcal{G}_{\bm{B}}=\{\bm{I}_{n}\}$ and the assumption is vacuous; $\mathsf{Col}(\bm{B})^{\perp}=\{0\}$ so our directional statements are degenerate. Conversely, $r=0$ reduces to ROI.
Appendix C Elliptical designs
This section presents the asymptotic normality and feature selection results for designs that violate the $\bm{B}$-right orthogonal invariance assumption, and provides numerical evidence that our theoretical results remain valid in more general settings.
As an instance of elliptical distributions, we examine a design where each row of $\bm{X}$ is independently drawn from a multivariate normal distribution with an AR(1) covariance structure. That is, $\bm{x}_{i}\overset{\rm iid}{\sim}\mathcal{N}(\bm{0},\bm{\Sigma})$ with $(\bm{\Sigma})_{ij}=\rho^{|i-j|}$, $\rho>0$ for any $i\in[n]$. We conducted experiments with $(m,n)=(2000,1000)$. To address potential instability induced by strong correlations, we set the learning rate and weight decay to $10^{-3}$ and $10^{-4}$, respectively, for the MLP and 1D-CNN, while keeping all other configurations identical to those in Section 5.1. Figure 4 presents the results, showing that the asymptotic normality of Theorem 1 for null variables is numerically preserved, irrespective of the correlation strength among features.
Figure 4. Histograms of the empirical distribution of ${\sqrt{n}\xi_{j}^{(10)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(10)}\|}$ for $j\in S^{\mathsf{c}}$. The solid black curve shows the $\mathcal{N}(0,1)$ density. The solid red curve represents a normal density fitted to the histograms.
Next, we examine the iteration-wise evolution of the FDR and Power observed during the numerical experiments. We fix the correlation parameter at $\rho=0.5$, set the learning rate and weight decay as described above, and keep all other settings identical to those in Section 5.2. The results are shown in Figure 5, indicating that FDR control is successfully achieved despite the presence of feature correlations. On the other hand, the detection power of LSTM begins to decay after a certain number of iterations, suggesting that early stopping could be beneficial.
Figure 5. Results for the FDR/power (left) and the training loss (right) when performing feature selection at each iteration. The solid curves represent averages over 20 independent runs, and the shaded areas indicate one standard deviation around the mean.
Appendix D Additional numerical experiments
This section provides additional and more detailed results complementing the experiments presented in Section 5.
QQ-plots. As further evidence supporting the asymptotic normality demonstrated in Section 5.1, Figure 6 shows the QQ-plots of ${\sqrt{n}\xi_{j}^{(10)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(10)}\|}$ for $j\in S^{\mathsf{c}}$.
Figure 6. QQ-plots of ${\sqrt{n}\xi_{j}^{(10)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(10)}\|}$ under the settings of Figure 2.
Loss trajectories. Figure 7 presents the evolution of the training loss corresponding to the FDR and Power trajectories shown in Figure 3.
Figure 7. Training loss trajectories of each method across iterations (corresponding to the methods plotted in Figure 3).
Classification problem. Finally, as an application to a different data-generating process, we consider a multi-class classification problem. Each entry of $\bm{X}\in\mathbb{R}^{m\times n}$ is drawn from an i.i.d. standard Gaussian distribution. We construct a weight matrix $\bm{B}\in\mathbb{R}^{n\times 3}$ by drawing a Gaussian matrix in $\mathbb{R}^{(n/2)\times 3}$, orthonormalizing its columns via QR, and embedding it into the top $n/2$ coordinates while filling zeros in $S^{\mathsf{c}}$. For $K=3$ classes, let the class-$k$ score be
$h_{k}(\bm{x})=\alpha_{k}\sin\big(\omega_{k}\,(\bm{B}_{\cdot 1}^{\top}\bm{x})\big)+\beta_{k}\cos\big(\nu_{k}\,(\bm{B}_{\cdot 2}^{\top}\bm{x})\big)+\gamma_{k}(\bm{B}_{\cdot 3}^{\top}\bm{x})+b_{k},$
where $(\alpha_{k},\beta_{k},\gamma_{k})$ control the relative contributions, $(\omega_{k},\nu_{k})$ set the frequencies and $b_{k}$ balances class prior. Class probabilities follow a softmax with temperature $\tau>0$,
$\Pr(y_{i}=k\mid\bm{x}_{i})=\frac{\exp\big(h_{k}(\bm{x}_{i})/\tau\big)}{\sum_{\ell=1}^{K}\exp\big(h_{\ell}(\bm{x}_{i})/\tau\big)},$
and labels are sampled accordingly. We keep $\tau$ and the amplitudes fixed across trials unless stated otherwise and vary the random seed to average over data realizations.
The configurations of the MLP, 1D-CNN, and LSTM models are the same as those used in Section 5. The choices of $m$ and $n$ follow the same setting as well. Figures 8 and 9 present histograms and QQ-plots that confirm the asymptotic normality of the proposed statistics. Figure 10 presents the results of feature selection when $m=4000$ and $n=400$. In this setting, the power remains nearly zero for all methods, and in some cases the training loss does not decrease. This behavior is likely due to the non-null distribution of $\xi_{j}^{(t)}$ not being sufficiently separated from its null counterpart, suggesting that further investigation is needed to determine whether this issue can be mitigated through adjustments to the network architecture or optimization strategy.
Figure 8. Histograms of the empirical distribution of ${\sqrt{n}\xi_{j}^{(10)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(10)}\|}$ for $j\in S^{\mathsf{c}}$ under the multi-class classification model. The solid black curve shows the $\mathcal{N}(0,1)$ density. The solid red curve represents a normal density fitted to the histograms, and the dotted blue line indicates the empirical mean.
Figure 9. QQ-plots of ${\sqrt{n}\xi_{j}^{(10)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(10)}\|}$ under the multi-class classification model.
Figure 10. Results for the FDR/power (left) and the training loss (right) when performing feature selection at each iteration under the multi-class classification model. The solid curves represent averages over 20 independent runs, and the shaded areas indicate one standard deviation around the mean.

### v2

@@
-Provable FDR Control for Deep Feature Selection: Arbitrarily Deep MLPs and Beyond
+Provable FDR Control for Deep Feature Selection: Deep MLPs and Beyond
 Abstract.
@@
 The method applies to architectures whose first layer is fully connected.
-From the second layer onward, it accommodates multilayer perceptrons of arbitrary width and depth, convolutional and recurrent networks, attention mechanisms, residual connections, and dropout.
+From the second layer onward, it accommodates multilayer perceptrons (MLPs) of arbitrary width and depth, convolutional and recurrent networks, attention mechanisms, residual connections, and dropout.
 The procedure also accommodates stochastic gradient descent with data-independent initializations and learning rates.
@@
 Two largely separate research threads have pursued this goal, namely high-dimensional statistics and explainable AI (XAI).
-On the statistics side, a central objective has been to provide theoretical guarantees on false discovery rate (FDR) control [abramovich2006adapting, meinshausen2006high, barber2015controlling, candes2018panning, xing2023controlling, dai2023false, dai2023scale, du2023false], often by debiasing sparse estimators like the LASSO [tibshirani1996regression].
+On the statistics side, a central objective has been to provide theoretical guarantees on false discovery rate (FDR) control [1, 15, 2, 6, 31, 9, 8, 10], often by debiasing sparse estimators like the LASSO [28].
 These guarantees, albeit rigorous, typically come at the expense of strong modeling assumptions (e.g., the true model follows a generalized linear model, or the feature distribution is known), which can diverge from complex real‑world phenomena and thereby cast doubt on the reliability of the selected features.
-On the XAI side, a wealth of attribution methods, e.g., LIME [ribeiro2016should], SHAP [lundberg2017unified], random forest feature importance [breiman2001random], and saliency maps [simonyan2013deep], quantify the contribution of individual features in black‑box models.
+On the XAI side, a wealth of attribution methods, e.g., LIME [19], SHAP [14], random forest feature importance [5], and saliency maps [24], quantify the contribution of individual features in black‑box models.
 Thresholding such scores yields a practical selection heuristic, yet without guarantees on Type‑I error; thus, error control has remained elusive.
@@
 In the former, e.g., for linear regression, whether coefficients are zero or nonzero determines which features should be selected, and FDR control can be achieved by invoking asymptotic normality of the estimators.
-Representative examples include the Gaussian mirror [xing2023controlling] and data splitting [dai2023false].
-A key advantage of these approaches is scale-free property, since they avoid estimating the asymptotic variance, and they have been extended to generalized linear models [dai2023scale].
-Similar techniques have also been applied to sliced inverse regression under multi-index models [zhao2022testing], although the theoretical assumptions on the feature dimension and the sparsity level are rather stringent.
-As the second line of research, the model-X knockoff [candes2018panning] is groundbreaking in that it can control the FDR without assuming the structure of $\bm{y}\mid\bm{X}$ where $\bm{y}\in\mathbb{R}^{m}$ is a response vector.
+Representative examples include the Gaussian mirror [31] and data splitting [9].
+A key advantage of these approaches is scale-free property, since they avoid estimating the asymptotic variance, and they have been extended to generalized linear models [8].
+Similar techniques have also been applied to sliced inverse regression under multi-index models [33], although the theoretical assumptions on the feature dimension and the sparsity level are rather stringent.
+As the second line of research, the model-X knockoff [6] is groundbreaking in that it can control the FDR without assuming the structure of $\bm{y}\mid\bm{X}$ where $\bm{y}\in\mathbb{R}^{m}$ is a response vector.
 However, it requires knowledge of the joint distribution of $\bm{X}$, which is restrictive.
-To address this limitation, methods that estimate the distribution of $\bm{X}$ using generative networks have been proposed [jordon2019knockoffgan, romano2020deep].
-In addition, a sequential knockoff sampler for a given feature distribution has been proposed [bates2021metropolized].
+To address this limitation, methods that estimate the distribution of $\bm{X}$ using generative networks have been proposed [11, 20].
+In addition, a sequential knockoff sampler for a given feature distribution has been proposed [3].
 1.1.2. FDR control via neural networks.
 Several studies have proposed FDR control methods that employ neural networks, although without theoretical guarantees and typically within restricted classes of architectures.
-The Neural Gaussian Mirror [xin2020ngm] defines a kernel-based conditional dependence measure and performs feature selection with MLPs.
-DeepPINK [lu2018deeppink] is a knockoff framework with a specially designed network architecture in which the features are assumed to be jointly Gaussian, and DeepLINK [zhu2021deeplink] relaxes this distributional restriction.
+The Neural Gaussian Mirror [30] defines a kernel-based conditional dependence measure and performs feature selection with MLPs.
+DeepPINK [13] is a knockoff framework with a specially designed network architecture in which the features are assumed to be jointly Gaussian, and DeepLINK [34] relaxes this distributional restriction.
 These approaches have been shown empirically to achieve FDR control.
@@
 Our goal is to select a relevant subset of feature indices from $[n]$ that are associated with $\bm{y}$.
-A desirable feature selection procedure controls the false discovery rate (FDR) [benjamini1995controlling], defined by
+A desirable feature selection procedure controls the false discovery rate (FDR) [4], defined by
 $\textsf{FDR}=\mathbb{E}[\textsf{FDP}],\quad\textsf{FDP}=\frac{\#\{j\notin S:j\in\hat{S}\}}{\#\{j:j\in\hat{S}\}\vee 1},$ (1)
@@
 Assumption 3 (Architecture of the neural network).
-The first layer of the network is taken to be dense and fully connected, and we denote its weight matrix by $\bm{W}_{1}\in\mathbb{R}^{n\times q}$.
+The first hidden layer of the network exists and is taken to be dense and fully connected, and we denote its weight matrix by $\bm{W}_{1}\in\mathbb{R}^{n\times q}$.
 The dependence of the entire network $f_{\mathcal{W}}(\bm{x})$ on any input $\bm{x}\in\mathbb{R}^{n}$ arises solely through the transformed representation $\bm{W}_{1}^{\top}\bm{x}$.
@@
 Proofs of the assertions in this section are deferred to Appendix A. Figure 2 exhibits that numerical results confirm the asymptotic normality of Theorem 1.
-The technical contribution underlying Proposition 1 and Theorem 1 is to extend the orthogonal-invariance-based theory of marginal asymptotic normality—originally developed by zhao2022asymptotic and later shown to be broadly applicable by sawaya2024high—from loss minimizers (i.e., M-estimators) to the individual iterates of loss-minimization algorithms.
+The technical contribution underlying Proposition 1 and Theorem 1 is to extend the orthogonal-invariance-based theory of marginal asymptotic normality—originally developed by Zhao et al.
+[32] and later shown to be broadly applicable by Sawaya et al.
+[21]—from loss minimizers (i.e., M-estimators) to the individual iterates of loss-minimization algorithms.
 As a consequence, our asymptotic normality results do not rely on
@@
 Points (ii) and (iii) are particularly important in the training of deep neural networks, where convergence to a global minimum is rarely guaranteed.
-Furthermore, our proofs extend naturally to gradient-based procedures for fitting conventional M-estimators without neural networks.
-This implies that valid inference remains possible even in settings where a maximum likelihood estimator may fail to exist—as in logistic regression [candes2020phase]—which further underscores the practical value of our approach.
+Example 3.1 (Linear regression).
+Among neural networks satisfying Assumption 3, the simplest example is the linear model $f_{\mathcal{W}}(\bm{x})=\bm{w}_{1}^{\top}\bm{x}$, i.e., single-layer neural networks.
+Since the multi-index model in Definition 1 includes the linear regression model $y_{i}=\bm{\beta}^{\top}\bm{x}_{i}+\varepsilon_{i}$ satisfying $\mathbb{E}[\varepsilon_{i}\mid\bm{x}_{i}]=0$, our result in Theorem 1 can be viewed as extending asymptotic normality results in high dimensions for arbitrary $M$-estimators satisfying Assumption 2 in linear regression to every iteration of stochastic gradient descent.
+Example 3.2 (Logistic regression).
+The multi-index model in Definition 1 also encompasses the logistic regression model $y_{i}=g(\bm{\beta}^{\top}\bm{x}_{i},\varepsilon_{i})$, where $g(u,v)=\mathbbm{1}\{(1+e^{-u})^{-1}>v\}$ and $\varepsilon_{i}\overset{\mathrm{iid}}{\sim}\mathrm{Unif}(0,1)$ independent of $\bm{x}_{i}$.
+By taking $f_{\mathcal{W}}(\bm{x})=\bm{w}_{1}^{\top}\bm{x}$ and the cross-entropy loss $\mathcal{L}(u,v)=\log(1+\exp(v))-uv$, our assumptions also cover the gradient descent iterates used to compute the maximum likelihood estimator.
+As shown by Candès and Sur [7], the maximum likelihood estimator in logistic regression may fail to exist.
+Notably, because our theory applies to each iterate of SGD, it remains valid even in such settings where the optimization trajectory diverges and convergence is not guaranteed.
 4. Methodology
@@
 According to Theorem 1, $\xi_{j}^{(t)}$ satisfies (a) but not (b), whereas $|\xi_{j}^{(t)}|$ satisfies (b) but not (a).
-Therefore, we adopt a data-splitting approach [dai2023false, dai2023scale].
+Therefore, we adopt a data-splitting approach [9, 8].
 Specifically, we randomly divide the data into two equal parts and, using the quantities $\xi_{j1}^{(t)}$ and $\xi_{j2}^{(t)}$ computed from each split, construct the importance statistics, for $j\in[n]$,
@@
 Such assumptions frequently appear in the related literature.
-Compared with Assumption 3.2 in dai2023scale, which requires that a fixed proportion of the true signals diverge, our assumption can be regarded as considerably weaker.
+Compared with Assumption 3.2 in Dai et al.
+[8], which requires that a fixed proportion of the true signals diverge, our assumption can be regarded as considerably weaker.
 Assumption 7 requires that, within the top $K_{n}$ statistics $M_{j}$’s ranked by magnitude, at least a fixed fraction corresponds to non-null and, moreover, lies on the positive side.
@@
 A drawback of the data-splitting method is that it effectively halves the sample size, which may reduce power.
-One remedy, known in the literature and also applicable here, is to aggregate the selection results obtained from multiple random splits of the data [dai2023false].
-Related stabilization techniques include du2023false and ren2024derandomised
+One remedy, known in the literature and also applicable here, is to aggregate the selection results obtained from multiple random splits of the data [9], as is commonly done in stability selection [16, 22].
+Related stabilization techniques include Du et al.
+[10] and Ren and Barber [18].
+Later, Takahashi [27] theoretically shows that such stabilization techniques increase power in the proportional asymptotics.
 5. Numerical experiments
@@
 We benchmark our approach against flexible variable-selection baselines that can (or aim to) control FDR.
-Neural baselines advertised as enabling FDR control via deep representations include the Neural Gaussian Mirror (NGM) [xin2020ngm] and DeepLINK [zhu2021deeplink].
+Neural baselines advertised as enabling FDR control via deep representations include the Neural Gaussian Mirror (NGM) [30] and DeepLINK [34].
 While empirically competitive, these methods do not offer formal guarantees.
 All methods are run at the same nominal FDR level $\alpha=0.1$.
-We ran DeepLINK [zhu2021deeplink] with authors’ official code [^1] https://github.com/zifanzhu/DeepLINK with hyperparameters as follows: L1 penalty $10^{-3}$, learning rate $10^{-3}$, ELU activation, and mean squared error loss.
+We ran DeepLINK [34] with authors’ official code [^1] https://github.com/zifanzhu/DeepLINK with hyperparameters as follows: L1 penalty $10^{-3}$, learning rate $10^{-3}$, ELU activation, and mean squared error loss.
 Column-wise centering and scaling were applied by default.
@@
 In this work, we considered the raw input gradients of a trained neural network as a feature importance.
-A promising direction for future work is to extend our analysis to more sophisticated feature attribution methods, including Integrated Gradients [sundararajan2017axiomatic], DeepLIFT [shrikumar2017learning], and SmoothGrad [smilkov2017smoothgrad].
+A promising direction for future work is to extend our analysis to more sophisticated feature attribution methods, including Integrated Gradients [26], DeepLIFT [23], and SmoothGrad [25].
 Adapting our FDR-control framework to such attribution methods would potentially yield more stable and interpretable feature selection.
@@
 Proof of Proposition 1.
-The basic idea of the proof is inspired by the proof of Proposition 2.1 in zhao2022asymptotic.
+The basic idea of the proof is inspired by the proof of Proposition 2.1 in Zhao et al.
+[32].
 Since $\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}/\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|$ has the unit norm and lies in $\mathrm{Col}(\bm{B})^{\perp}$, it is sufficient to show that, for any orthogonal matrix $\bm{U}\in\mathbb{R}^{n\times n}$ obeying $\bm{U}\bm{B}=\bm{B}$,
@@
 Since $\mathsf{rank}(\bm{B})=q^{*}$ by Assumption 1 (i), we have $\|\bm{P}_{\bm{B}}^{\perp}\|_{F}^{2}=n-q^{*}$ and $\|\bm{P}_{\bm{B}}^{\perp}\|_{\rm op}=1$.
-Thus, Hanson–Wright inequality [vershynin2018high] implies that, for any $t>0$,
+Thus, Hanson–Wright inequality [29] implies that, for any $t>0$,
 $\mathbb{P}\left(\left\lvert\bm{z}^{\top}\bm{P}_{\bm{B}}\bm{z}-(n-q^{*})\right\rvert>t\right)\leq 2\exp\left(-c\min\left(\frac{t^{2}}{n-q^{*}}\wedge t\right)\right),$ (32)
@@
 Also, since $\sigma_{n}\in[1-\epsilon,1+\epsilon]$ under $E_{n}$, it follows that
-$\mathbb{P}\left(z_{j}<\frac{u}{1+\epsilon}\right)\leq\mathbb{P}\left(\sigma_{n}z_{j}\mid E_{n}\right)\leq\mathbb{P}\left(z_{j}<\frac{u}{1-\epsilon}\right).$ (40)
+$\mathbb{P}\left(z_{j}<\frac{u}{1+\epsilon}\right)\leq\mathbb{P}\left(\sigma_{n}z_{j}<u\mid E_{n}\right)\leq\mathbb{P}\left(z_{j}<\frac{u}{1-\epsilon}\right).$ (40)
 From this, we have
@@
 A.3. Proof of Thorem 2
-Our argument is inspired by the proof of Proposition 3.2 of dai2023scale, but proceeds under weaker conditions on signal strength and dimensionality.
-While auxiliary lemmas overlap with dai2023scale, we provide full proofs to ensure a self-contained presentation.
+Our argument is inspired by the proof of Proposition 3.2 of Dai et al.
+[8], but proceeds under weaker conditions on signal strength and dimensionality.
+While auxiliary lemmas overlap with Dai et al.
+[8], we provide full proofs to ensure a self-contained presentation.
 We prepare some notation for the proof.
@@
 Lemma A.5.
+Under the assumptions of Theorem 1, as $n\to\infty$ while $q^{*}=o(n)$, we have,
+$\sup_{i,j\in S^{\mathsf{c}},\,t_{1},t_{2}\in\mathbb{R}}\left\lvert\mathbb{P}\left(\frac{\sqrt{n}\xi_{i}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}<t_{1},\frac{\sqrt{n}\xi_{j}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}<t_{2}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert\to 0.$ (58)
+Proof of Lemma A.5.
+We assume, without loss of generality, that $t_{1}>0$ and $t_{2}>0$.
+Using the fact that ${\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}\overset{\rm d}{=}{\bm{P}_{\bm{B}}^{\perp}\bm{z}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{z}\|}$ where $\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$ by Theorem 1, we have,
+$\left(\frac{\sqrt{n}\xi_{i}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|},\frac{\sqrt{n}\xi_{j}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}\right)\overset{\rm d}{=}\left(\frac{\sqrt{n}z_{i}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{z}\|},\frac{\sqrt{n}z_{j}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{z}\|}\right),$ (59)
+for $i,j\in S^{\mathsf{c}}$.
+This follows from the fact that the $j$-th element of ${\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}$ is ${\xi_{j}^{(t)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}$ under the null since $\bm{b}_{j}=\bm{0}$ for $j\in S^{\mathsf{c}}$.
+Denote $\sigma_{n}=\sqrt{n}/\|\bm{P}_{\bm{B}}^{\perp}\bm{z}\|$.
+From (59), it follows that
+$\Delta_{n}$ $\equiv\sup_{i,j\in S^{\mathsf{c}},\,t_{1},t_{2}\in\mathbb{R}}\left\lvert\mathbb{P}\left(\frac{\sqrt{n}\xi_{i}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}<t_{1},\frac{\sqrt{n}\xi_{j}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}<t_{2}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert$ (60)
+$=\sup_{i,j\in S^{\mathsf{c}},\,t_{1},t_{2}\in\mathbb{R}}\left\lvert\mathbb{P}\left(z_{i}<\sigma_{n}t_{1},z_{j}<\sigma_{n}t_{2}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert$. (61)
+Fix an arbitrary $\epsilon>0$ and denote $E_{n}=\{|\sigma_{n}-1|<\epsilon\}$. Then we have
+$\left\lvert\mathbb{P}\left(z_{i}<\sigma_{n}t_{1},z_{j}<\sigma_{n}t_{2}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert$ (62)
+$\leq\mathbb{P}(E_{n}^{\mathsf{c}})+\left\lvert\mathbb{P}\left(z_{i}<\sigma_{n}t_{1},z_{j}<\sigma_{n}t_{2}\mid E_{n}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert$. (63)
+Also, since $\sigma_{n}\in[1-\epsilon,1+\epsilon]$ under $E_{n}$, it follows that
+$\mathbb{P}\left(z_{i}<\frac{t_{1}}{1+\epsilon},z_{j}<\frac{t_{2}}{1+\epsilon}\right)\leq\mathbb{P}\left(z_{i}<\sigma_{n}t_{1},z_{j}<\sigma_{n}t_{2}\mid E_{n}\right)\leq\mathbb{P}\left(z_{i}<\frac{t_{1}}{1-\epsilon},z_{j}<\frac{t_{2}}{1-\epsilon}\right).$ (64)
+From this, we have
+$\left\lvert\mathbb{P}\left(z_{i}<\sigma_{n}t_{1},z_{j}<\sigma_{n}t_{2}\mid E_{n}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert$ (65)
+$\leq\max\left\{\Phi\left(\frac{t_{1}}{1-\epsilon}\right)\Phi\left(\frac{t_{2}}{1-\epsilon}\right)-\Phi(t_{1})\Phi(t_{2}),\,\Phi(t_{1})\Phi(t_{2})-\Phi\left(\frac{t_{1}}{1+\epsilon}\right)\Phi\left(\frac{t_{2}}{1+\epsilon}\right)\right\}.$ (66)
+Taking the supremum, the mean-value theorem gives
+$\sup_{t_{1},t_{2}\in\mathbb{R}}\left\lvert\mathbb{P}\left(z_{i}<\sigma_{n}t_{1},z_{j}<\sigma_{n}t_{2}\mid E_{n}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert\leq\frac{\epsilon}{1-\epsilon}\cdot\sqrt{\frac{2}{\pi e}},$ (67)
+where we use the fact $\sup_{u\in\mathbb{R}}|u|\phi(u)=1/\sqrt{2\pi e}$. Since this upper bound does not depend on $j\in[n]$, (63) and (67) yield
+$\Delta_{n}\leq\mathbb{P}(|\sigma_{n}-1|>\epsilon)+\frac{\epsilon}{1-\epsilon}\cdot\sqrt{\frac{2}{\pi e}}.$ (68)
+The first term on the right-hand side converges to zero as $n\to\infty$ while $q^{*}=o(n)$ by Lemma A.1, and the second term goes to zero as $\epsilon\downarrow 0$. ∎
+Lemma A.6.
 Let $n_{0}$ be the number of null features. Under the assumptions of Theorem 2, as $n\to\infty$ while $q^{*}=o(n)$, we have
@@
 $\leq\mathbb{E}\left[\mathbb{P}(z_{j^{\prime}2}>I_{u}(x),\,z_{j2}>I_{u}(y))\mid\varsigma_{j^{\prime}1}^{(t)}=x>0,\,\varsigma_{j1}^{(t)}=y>0\right]+2\Delta$ (82)
-$=\mathbb{E}\left[Q(I_{u}(x))Q(I_{u}(y))\mid\varsigma_{j^{\prime}1}^{(t)}=x>0,\,\varsigma_{j1}^{(t)}=y>0\right]+2\Delta.$ (72)
-Similarly, we can upper bound $I_{2}$, $I_{3}$, and $I_{4}$. Combining the four upper bounds together, we obtain an upper bound on $\mathbb{P}(\breve{M}_{j}>u,\breve{M}_{j^{\prime}}>u)$ as
+$=\mathbb{E}\left[Q(I_{u}(x))Q(I_{u}(y))\mid\varsigma_{j^{\prime}1}^{(t)}=x>0,\,\varsigma_{j1}^{(t)}=y>0\right]+2\Delta,$ (83)
+where the inequality follows from Lemma A.5.
+Similarly, we can upper bound $I_{2}$, $I_{3}$, and $I_{4}$.
+Combining the four upper bounds together, we obtain an upper bound on $\mathbb{P}(\breve{M}_{j}>u,\breve{M}_{j^{\prime}}>u)$ as
 $\mathbb{P}\left(\mathsf{sign}(z_{j2}\varsigma_{j1}^{(t)})\psi(|z_{j2}|,|\varsigma_{j1}^{(t)}|)>u,\ \mathsf{sign}(z_{j^{\prime}2}\varsigma_{j^{\prime}1}^{(t)})\psi(|z_{j^{\prime}2}|,|\varsigma_{j^{\prime}1}^{(t)}|)>u\right)+8\Delta.$ (84)
@@
 In this section, we delineate what kinds of random designs fall into the class of right-orthogonally invariant (ROI) matrices, and what kinds do not.
-ROI is sometimes assumed in the literature of approximate message passing algorithms in the proportional asymptotics where $n$ and $m$ diverge with $m/n\to\delta\in(0,\infty)$ [rangan2019vector, li2023spectrum].
+ROI is sometimes assumed in the literature of approximate message passing algorithms in the proportional asymptotics where $n$ and $m$ diverge with $m/n\to\delta\in(0,\infty)$ [17, 12].
 Subsequently, we discuss in what sense the $\bm{B}$-ROI in Assumption 1 (ii) is relaxed.
@@
 i.e. the conditional law of $\bm{X}\bm{P}_{\bm{B}}^{\perp}$ given $\bm{X}\bm{P}_{\bm{B}}$ is ROI.
-New important examples under $\bm{B}$-ROI.
+New important examples under $\bm{B}$-ROI are
   • Fixed-loading spike + isotropic noise: $\bm{X}=\bm{F}\bm{\Lambda}^{\top}+\bm{E}$ with $\mathsf{Col}(\bm{\Lambda})=\mathsf{Col}(\bm{B})$ and $\bm{E}$ isotropic on $\mathsf{Col}(\bm{B})^{\perp}$.